### v1

# Dynamic Attention-Guided Context Decoding for Mitigating Context Faithfulness Hallucinations in Large Language Models

Abstract
Large language models (LLMs) often suffer from context faithfulness hallucinations, where outputs deviate from retrieved information due to insufficient context utilization and high output uncertainty. Our uncertainty evaluation experiments reveal a strong correlation between high uncertainty and hallucinations. We hypothesize that attention mechanisms encode signals indicative of contextual utilization, validated through probing analysis. Based on these insights, we propose Dynamic Attention-Guided Context Decoding (DAGCD), a lightweight framework that integrates attention distributions and uncertainty signals in a single-pass decoding process. Experiments across QA datasets demonstrate DAGCD’s effectiveness, achieving significant improvements in faithfulness and robustness while maintaining computational efficiency.
Dynamic Attention-Guided Context Decoding for Mitigating Context Faithfulness Hallucinations in Large Language Models
Yanwen Huang1,2,†, Yong Zhang1,†, Ning Cheng1,*, Zhitao Li1, Shaojun Wang1, Jing Xiao1, 1 Ping An Technology (Shenzhen) Co., Ltd., China 2 University of Electronic Science and Technology of China {zhangyong203, chengning211}@pingan.com.cn
footnotetext: † Equal contribution.footnotetext: * Corresponding author.footnotetext: This work was done during Yanwen Huang’s internship at Ping An Technology (Shenzhen) Co., Ltd., China.
1 Introduction
Large Language Models (LLMs) Brown et al. (2020); Achiam et al. (2023); Touvron et al. (2023) excel at generating fluent and contextually relevant responses. However, they often struggle with producing factually accurate outputs, particularly when relying on external information Vu et al. (2023). Retrieval-Augmented Generation (RAG) Lewis et al. (2020); Guu et al. (2020) mitigates this by grounding outputs in retrieved context, making it effective for tasks like question answering and reasoning Gao et al. (2023); Fan et al. (2024). However, models often fail to faithfully utilize retrieved context, resulting in context faithfulness hallucinations, where outputs deviate from the retrieved context Huang et al. (2023a); Ji et al. (2023). These issues undermine RAG systems’ reliability, particularly in high-stakes applications requiring factual accuracy Chuang et al. (2024a).
Figure 1: Analysis of generation distribution when concatenating context: ($a_{1},a_{2}$) the model’s uncertainty when generating correct versus wrong answers; ($b$) for wrong answers, the ranking of the gold answer token in the distribution, and the probability difference between the gold answer token and the Top-1 predicted token.
Existing methods, such as CAD Shi et al. (2024b) and COIECD Yuan et al. (2024), attempt to mitigate context faithfulness hallucinations by dynamically adjusting decoding distributions through generation distribution comparisons or token-level uncertainty signals. While effective to some extent, these methods face several key limitations: limited interpretability, degraded performance when context-agnostic and context-aware outputs differ significantly, and high computational overhead due to multiple decoding passes. These issues highlight the need for lightweight and interpretable approaches that can effectively utilize retrieved context while maintaining scalability.
To further explore these issues, we analyze the relationship between output uncertainty and context faithfulness hallucinations. As shown in Figure 1, we observe a strong correlation between context hallucinations and high entropy in the generation distribution. Specifically, incorrect generations often reflect significant uncertainty through high entropy. Notably, in cases of wrong answers, context tokens corresponding to the correct answer often appear within the top-10 ranks but are assigned insufficient confidence. This indicates that the model’s internal mechanisms encode interpretable signals to indicate context utilization.
Building on our analysis and findings in interpretability studies, we hypothesize that attention mechanisms in Transformer models encode signals indicative of contextual utilization. Attention modules play a critical role in aligning generated tokens with their corresponding context tokens by distributing focus across input tokens Geva et al. (2023); Jin et al. (2024b). To validate our hypothesis, we trained a probing classifier using Logistic Regression on attention distributions, achieving over 97% accuracy in distinguishing contextually relevant tokens. These results demonstrate that attention weights can serve as a lightweight and interpretable proxy for assessing the effective incorporation of retrieved evidence during generation.
Motivated by these findings, we propose Dynamic Attention-Guided Context Decoding (DAGCD), a novel algorithm to mitigate context faithfulness hallucinations. Inspired by the copy-generator framework See et al. (2017); Xu et al. (2020b), DAGCD integrates attention weights to estimate the relevance of tokens in the retrieved context, dynamically adjusting output probabilities. Entropy-based uncertainty measures further refine these adjustments, amplifying underconfident but contextually relevant tokens. By unifying these strategies, DAGCD aligns outputs with context while maintaining computational efficiency.
Our contributions are as follows:
  1. 1.
    Comprehensive analysis of context faithfulness hallucinations: We reveal a strong correlation between context faithfulness hallucinations and output uncertainty through entropy-based analysis, revealing the model’s challenges in aligning with retrieved context under high uncertainty.
  2. 2.
    Attention-driven interpretability framework: We propose Dynamic Attention-Guided Context Decoding (DAGCD), leveraging attention distributions to amplify contextually relevant tokens and ensure faithful utilization of retrieved context.
  3. 3.
    Lightweight and efficient decoding: DAGCD operates in a single decoding pass, integrating attention signals and uncertainty measures without additional overhead, achieving improved decoding efficiency.
  4. 4.
    Extensive validation across datasets and models: DAGCD outperforms greedy decoding across multiple QA datasets, achieving an average F1 score improvement of 8.95% on pretrained models and 1.91% on instruction-tuned models, demonstrating its robustness and scalability.
2 Why Can’t Generate Faithful Answers?
LLMs often fail to generate responses that are faithful to the concatenated context. We analyzed (1) the impact of context concatenation on output uncertainty (§2.1) and (2) the ranking of contextual answer tokens within predictions (§2.2). Our findings indicate that high output uncertainty and low confidence in contextually relevant tokens are critical factors contributing to this issue.
2.1 Uncertainty Leads to Unfaithful Answers
To assess the relationship between uncertainty and response faithfulness, we focus on two widely used metrics: entropy, which measures the overall uncertainty in the output distribution Xu et al. (2020a); van der Poel et al. (2022); Huang et al. (2023b), and top-1 token probability, which reflects the model’s confidence in its highest-ranked prediction Yao et al. (2023); Luo et al. (2024). Higher entropy indicates greater uncertainty, while higher top-1 probability corresponds to greater confidence in the model’s predictions.
Experimental Setup We compute the normalized entropy and top-1 token probability for the first generated token. Normalized entropy is defined as:
$H_{\text{norm}}(P)=-\frac{\sum_{i=1}^{N}P_{i}\log P_{i}}{\log N},$ (1)
where $P$ represents the probability distribution, and $N$ denotes the vocabulary size. These metrics quantify prediction confidence, with lower entropy and higher top-1 probability reflecting greater certainty. Detailed descriptions of the datasets, models, and calculation processes are provided in Appendix A.
Results and Analysis Figure 1 (subplots $a_{1}$ and $a_{2}$) illustrates the normalized entropy and top-1 token probability distributions for correct and wrong answers. The model exhibits higher uncertainty for wrong answers, with an average normalized entropy of 0.36 compared to 0.29 for correct cases, and a lower average top-1 probability of 0.25 compared to 0.41. Notably, a significant overlap exists between the entropy distributions of correct (blue) and wrong (red) cases, indicating that even correct answers often exhibit high uncertainty.
Model w/o context w/ context
LLaMA2-7B -0.43 -0.53
LLaMA2-7B-Chat -0.27 -0.33
LLaMA2-13B -0.30 -0.51
LLaMA2-13B-Chat -0.26 -0.32
Mistral-7B -0.23 -0.22
Mistral-7B-Instruct -0.30 -0.33
Table 1: Spearman correlation analysis. We examine the relationship between F1 scores (pred-ans vs. gold-ans) and norm-entropy (generation distribution) under both w/o and w/ context settings ($p\ll 0.05$ for all models).
Building on this observation, we conduct a Spearman correlation analysis (Table 1), further demonstrating how uncertainty (norm-entropy of the generation distribution) negatively correlates with accuracy (F1 score between predicted and golden answers) across all models. After concatenating context, this correlation becomes even more pronounced, confirming that uncertainty is a key factor in generating unfaithful answers.
2.2 LLM is Actually Utilizing Context
Experimental Setup Using the wrong answer samples from §2.1, we analyze the ranking distribution of the gold answer token in the first generated distribution following context concatenation. Ranks are grouped into intervals (e.g., rank 2–4, rank 5–7, etc.), and we compute the proportion of gold answer tokens within each rank interval. Additionally, we calculate the average probability gap between the gold answer token and the top-1 token.
Results and Analysis As shown in Figure 1 (subplot $b$), when the model answers incorrectly after context concatenation, 43% of the samples have the gold answer token ranked between 2–4, and a total of 66% have the gold answer token ranked within the top 10. Moreover, the average probability gap between the gold answer token and the top-1 token is relatively small: only 0.14 for ranks between 2–4 and 0.24 for ranks greater than 30. This indicates that, in most cases, the model can correctly identify the gold answer token within the context after concatenation, but it often fails to prioritize it effectively during the generation process.
The findings highlight that in context faithfulness hallucination scenarios, the model utilizes correct context tokens but assigns them insufficient confidence, limiting their influence on the output. This limitation underscores the model’s incomplete integration of retrieved context, prompting an exploration of how context tokens are incorporated to guide strategies for enhancing faithfulness and mitigating hallucinations.
3 Context Utilization Signal in Attention
3.1 Background and Hypothesis
Building on our findings, we focus on the attention mechanisms in Transformer models as a potential indicator of context token utilization. Attention mechanisms distribute focus across input tokens during generation, capturing the alignment between generated tokens and their corresponding context tokens Clark et al. (2019a); Geva et al. (2023). We hypothesize that attention heads encode signals indicating the contribution of retrieved context tokens to the generation of specific outputs. This provides a lightweight and interpretable approach to understanding contextual integration.
3.2 Attention Ratio
Analyzing raw attention weights poses challenges, as attention sink often results in disproportionate focus on non-context tokens such as delimiters, introducing noise to the analysis.Xiao et al. (2024); Clark et al. (2019b); Bondarenko et al. (2021); Son et al. (2024); Sun et al. (2024). Additionally, varying attention magnitudes across heads and layers Jain and Wallace (2019); Vig and Belinkov (2019) complicate feature comparison. To address these challenges, we compute normalized attention ratios exclusively for context tokens, reducing noise and stabilizing values across layers and heads. This approach constrains features to $[0,1]$, enabling consistent evaluation of context token importance. Formally, the attention ratio for token $j$ during the generation step, at the $h$-th head in the $l$-th layer, is defined as:
$r_{l,h}^{j}=\frac{a_{l,h}^{j}}{\sum_{j\in C}a_{l,h}^{j}}$ (2)
where $a_{l,h}^{j}$ represents the raw attention weight assigned to token $j$ in the context $C$, and $\sum_{j\in C}a_{l,h}^{j}$ is the total attention weight for all tokens in $C$. This ratio quantifies the relative importance of token $j$ within the context for a specific attention head.
For each token $j$, we construct a feature vector by aggregating its attention ratios across all heads:
$v_{j}=\left[r_{1,1}^{j},\dots,r_{\text{num\_layers},\text{num\_heads}}^{j}\right]$ (3)
the attention ratio feature vector $v_{j}$ provides a comprehensive representation of the token’s attention patterns, capturing its relative importance across the model’s entire architecture.
3.3 Experimental Setup and Results
To validate our hypothesis, we focused on cases where the model’s output shifted from incorrect to correct after context concatenation. Context tokens were labeled as positive (utilized) if they corresponded to the gold answer, and negative (non-utilized) otherwise. Using these labels, we extracted attention ratio feature vectors $\mathbf{v}_{j}$ under Prompt 1 to train a Logistic Regression (LR) classifier. The classifier was evaluated across multiple Transformer-based LLMs, including LLaMA2 and Mistral models. Details on dataset construction, training, and evaluation are in Appendix B.
Model ACC AUC
LLaMA2-7B 0.9793 0.9925
LLaMA2-7B-Chat 0.9798 0.9970
LLaMA2-13B 0.9740 0.9935
LLaMA2-13B-Chat 0.9737 0.9924
Mistral-7B 0.9787 0.9915
Mistral-7B-Instruct 0.9756 0.9883
Table 2: Performance of the LR classifier. We report the average performance across different random seeds.
The LR classifier demonstrates excellent performance across all evaluated LLMs, achieving accuracy above 0.97 and AUC scores exceeding 0.99 (Table 2). These results confirm that attention ratio features effectively distinguish utilized tokens from non-utilized ones, validating the hypothesis that attention patterns encode signals of contextual utilization. This robust performance, observed across various LLaMA2 and Mistral models, including chat and instruction-tuned variants, highlights the generality of attention ratio features across different architectures and fine-tuning paradigms. Moreover, the findings emphasize the potential of attention-based features as lightweight, interpretable signals for improving RAG systems.
3.4 Generalization and Feature Analysis
To validate the robustness and generalization of the LR classifier, we conducted extensive evaluations across out-of-domain data, prompt variations, and feature contributions.
Out-of-Domain Evaluation We tested the classifier on six sub-datasets from different domains, training on five sub-datasets and evaluating on the held-out one. As shown in Figure 2, the classifier achieves an average accuracy (ACC) and area under the curve (AUC) above 0.95 across all datasets and LLMs. These results demonstrate strong generalization capabilities of attention ratio features, even when applied to unseen domains.
Additional Analyses We further evaluated the robustness of attention-based features to prompt variations and analyzed the importance and characteristics of individual features. Results show that the classifier consistently performs well across prompts with minimal variation and the performance of classifier with top-K features surpasses that of using full features. Additionally, attention heads exhibit complementary behavior and data-dependent characteristics. Detailed results for these analyses are provided in Appendix C and Appendix D.
Figure 2: Out-of-domain performance of the LR classifier. One dataset is used as the test set, and the remaining datasets are used as the training set.
Figure 3: The illustration of the generation process of our proposed DAGCD method.
4 Method
Inspired by the copy-generation mechanism See et al. (2017); Xu et al. (2020b), we propose Dynamic Attention-Guided Context Decoding (DAGCD) to migrate context faithfulness hallucinations. DAGCD leverages utilization signals to dynamically guide the generation process, focusing on relevant contextual tokens. It integrates three steps: detecting utilized context tokens during inference (§4.1), constructing a utilization distribution (§4.2), and adjusting the generation probabilities to enhance contextual relevance (§4.3).
4.1 Context Utilization Detection
Context Utilization Detector In Section 3.4, we observed that the LR classifier trained using attention ratios from all attention heads did not generalize as effectively as the LR classifier trained with top-K features. Consequently, we selected the LR model trained with top-K features as the Context Utilization Detector. The set of attention heads corresponding to the top-K features used for training is denoted as $A=[A_{l_{1},h_{1}},A_{l_{2},h_{2}},\dots,A_{l_{K},h_{K}}]$, where each head $A_{l_{k},h_{k}}$ is associated with a feature coefficient $c_{k}$ (for all models, we set $K$ = 10).
[…]
$\mathbf{v}_{j}^{(K)}=[r_{l_{1},h_{1}}^{j},r_{l_{2},h_{2}}^{j},\dots,r_{l_{K},h_{K}}^{j}]$ (4)
Finally, the feature vector $\mathbf{v}_{j}^{(K)}$ is fed into the detector, which identifies the set of context tokens actively utilized at the current decoding step.
4.2 Utilization Distribution Construction
The context utilization detector identifies utilized tokens but does not quantify the degree of utilization for each token. To address this, we aggregate the attention ratios from different attention heads, weighted by their normalized feature coefficients $w_{k}$, to compute a utilization score for each token. Formally, the utilization score $s_{j}$ for token $j$ is defined as:
$s_{j}=\sum_{k=1}^{K}(r_{l_{k},h_{k}}^{j}\times w_{k}),\ w_{k}=\frac{c_{k}}{\sum_{k=1}^{K}c_{k}}$ (5)
The feature coefficient reflects the significance of each attention head, with positive values indicating a higher likelihood of token utilization and negative values suggesting a reduced likelihood. The utilization distribution is then constructed based on the utilization scores of individual tokens:
$\mathbf{U}=[u_{1},u_{2},\dots,u_{N}],\quad u_{i}=\frac{s_{i}}{\sum_{j=1}^{N}s_{j}}$ (6)
where $u_{i}$ represents the normalized utilization score for token $i$. Here, $N$ denotes the size of the vocabulary, and $s_{i}=0$ for tokens either absent from the context or not utilized, resulting in $u_{i}=0$ for such tokens.
4.3 Generate More Faithful Answers
Building on the utilization signals and the utilization distribution, DAGCD refines the generation process by dynamically adjusting token probabilities to prioritize contextually relevant tokens.
[…]
Dynamic Scaling of Adjustments The intensity of adjustments is guided by the model’s output uncertainty, measured as the normalized entropy $H_{\text{norm}}(P)$ of the generation distribution, as defined in Equation 1. High entropy indicates greater uncertainty and a higher risk of generating contextually inconsistent responses. Since our analysis shows that entropy correlates with uncertainty but does not establish a direct numerical relationship, we introduce a scaling factor $\alpha$, which compensates for model-specific entropy variations.
Top-Rank Filtering and Distribution Adjustment To enhance the reliability of generation adjustments, our approach incorporates a top-rank restriction, ensuring modifications focus on plausible tokens. Specifically, we define $U_{\text{top}}$ as the subset of the utilization distribution $U$ corresponding to tokens ranked within the top-$R$ positions of the generation distribution. This design addresses challenges similar to those highlighted in prior work Li et al. (2023); Chuang et al. (2024b), while specifically leveraging the observation that correct context tokens are typically found within the top-rank set, even when the model produces incorrect answers. Restricting adjustments to $U_{\text{top}}$ reduces the risk of amplifying irrelevant or nonsensical tokens, thereby maintaining the output’s integrity.
Under this framework, adjustments are applied only when utilized tokens in $U_{\text{top}}$ overlap with the top-rank set of the generation distribution. The adjusted generation distribution $P^{\prime}$ is computed as:
$P^{\prime}=P+\alpha H_{\text{norm}}(P)\cdot U_{\text{top}}$ (7)
where $P$ is the original generation distribution, and $\alpha H_{\text{norm}}(P)$ scales the adjustment intensity dynamically based on the normalized entropy $H_{\text{norm}}(P)$ of the entire generation distribution.
5 Experiments
Dataset Decoding HotpotQA TriviaQA SearchQA SQuAD NewsQA NQ NQ-swap Average
EM F1 EM F1 EM F1 EM F1 EM F1 EM F1 EM F1 EM F1
LLaMA2-7B Regular 44.74 54.71 55.28 68.01 54.21 59.24 39.90 52.61 32.93 45.46 38.85 50.59 36.03 36.62 43.13 52.46
CAD 44.13 54.49 55.26 68.04 54.14 59.21 38.35 51.12 31.74 43.70 38.14 48.58 36.11 36.69 42.55 51.69
COIECD 42.03 51.48 57.04 70.06 57.03 63.14 40.93 54.78 34.40 48.48 38.79 51.69 34.98 35.63 43.60 53.61
OURs 47.67 57.87 58.46 71.42 55.04 60.35 49.00 61.04 36.70 49.29 48.10 60.71 54.33 54.95 49.90 59.38
LLaMA2-7B -Chat Regular 53.33 67.41 71.72 76.83 54.19 58.11 67.69 78.70 39.41 53.90 50.47 65.48 67.98 68.78 57.83 67.03
CAD 52.86 67.16 71.70 76.83 54.16 58.11 65.89 77.91 38.46 53.26 48.89 65.00 68.04 68.85 57.14 66.73
COIECD 53.14 67.03 72.26 77.33 55.04 58.98 68.32 79.56 40.00 54.55 52.39 66.84 69.48 70.13 58.66 67.77
OURs 55.26 68.65 70.05 76.56 54.70 58.57 68.60 79.65 40.38 54.72 51.75 66.93 69.50 70.30 58.61 67.91
LLaMA2-13B Regular 52.36 63.40 58.25 69.95 63.22 68.33 51.64 64.57 30.84 40.11 42.26 54.08 49.02 49.59 49.66 58.58
CAD 51.53 63.11 58.25 69.92 63.13 68.32 49.94 63.44 29.94 39.04 41.40 52.27 49.07 49.63 49.04 57.96
COIECD 50.21 60.96 59.19 71.49 65.97 71.00 52.73 65.93 35.66 50.58 42.35 54.37 48.35 48.84 50.64 60.45
OURs 53.86 65.25 59.65 72.07 61.41 67.08 57.36 69.85 33.52 42.81 55.34 71.23 65.86 66.19 55.29 64.93
LLaMA2-13B -Chat Regular 55.01 69.92 74.58 79.35 67.08 71.96 68.26 79.45 40.20 55.11 53.49 69.18 60.69 61.77 59.90 69.53
CAD 54.44 69.66 74.58 79.37 67.01 71.95 66.77 78.70 39.44 54.44 52.89 68.63 60.83 61.92 59.42 69.24
COIECD 56.15 70.43 73.87 78.96 67.28 71.93 68.49 80.39 40.75 56.16 53.69 69.81 62.47 63.21 60.39 70.13
OURs 58.22 71.99 73.55 79.11 68.19 72.73 69.83 88.88 40.77 56.54 55.34 71.24 64.30 65.42 61.46 72.27
Mistral-7B Regular 53.41 64.36 59.45 68.39 63.79 67.77 44.19 56.11 31.51 38.94 33.74 51.18 39.80 46.04 46.56 56.11
CAD 41.57 56.01 57.88 67.48 63.64 68.65 34.08 47.55 25.78 35.37 23.18 41.46 26.96 35.97 39.01 50.36
COIECD 46.43 58.32 44.30 51.77 54.82 59.17 50.50 60.98 40.05 52.78 42.12 56.58 59.53 61.89 48.25 57.36
OURs 63.43 73.41 57.10 72.01 64.71 70.62 62.44 73.86 41.41 55.00 57.80 72.54 69.90 70.39 59.54 69.69
Mistral-7B -Instruct Regular 58.70 72.18 69.64 75.61 44.42 49.63 67.28 79.37 39.79 54.72 52.29 66.93 66.90 67.83 57.00 66.61
CAD 49.30 64.81 70.23 75.95 45.42 50.96 59.97 72.92 34.97 51.90 42.63 58.49 52.04 53.75 50.65 61.25
COIECD 59.74 72.59 64.92 72.15 37.09 42.66 68.45 81.03 40.84 55.96 53.54 68.72 72.81 73.81 56.77 66.70
OURs 60.77 73.82 69.70 75.96 47.17 52.65 68.29 80.74 40.85 56.01 54.69 69.81 71.27 72.10 58.96 68.73
Table 3: Performance comparison of different decoding methods. All baselines are reproduced under the same settings. Bold indicates the best performance, and underlined indicates the second-best performance.
5.1 Experimental Setup
Models We conducted experiments on LLaMA and Mistral, including LLaMA2 (7B, 13B, 7B-Chat, 13B-chat) Touvron et al. (2023) and Mistral (7B, 7B-Instruct) Jiang et al. (2023).
Datasets We conducted experiments on seven open-book question-answering (QA) datasets, representing a variety of QA tasks. These include multi-hop reasoning datasets (HotpotQA Yang et al. (2018)), long-form retrieval-based QA datasets (TriviaQA Joshi et al. (2017) and SearchQA Dunn et al. (2017)), single-paragraph extraction tasks (SQuAD Rajpurkar et al. (2016) and NewsQA Trischler et al. (2017)), and document-level QA datasets (NQ Kwiatkowski et al. (2019)). All datasets are formatted in the unified schema provided by the MrQA repository Fisch et al. (2019). Additionally, we used the artificially constructed NQ-swap dataset Longpre et al. (2021), designed to simulate conflicting or ambiguous scenarios by replacing entities. Detailed descriptions see Appendix F.3.
Baselines We compare the proposed DAGCD method with three decoding strategies: Greedy Decoding, CAD Shi et al. (2024b), and COIECD Yuan et al. (2024). CAD and COIECD are specifically designed to mitigate context faithfulness hallucination.
Implementation Details For DAGCD, the scaling factor $\alpha$ is set to 2 for pretrained models and 4 for instruction-tuned models to account for entropy variations. The logistic regression classifier and utilization distribution are computed using the top-10 attention heads, with adjustments restricted to the top-10 ranked tokens ($U_{\text{top}}$). Further details on prompts, decoding strategies, and baseline configurations are provided in Appendix F.4 and F.5.
Metrics Consistent with prior workJin et al. (2024a); Yuan et al. (2024); Wang et al. (2024), we use EM and F1 score metrics to evaluate the performance of the models on open-domain QA datasets.
5.2 Model Performance Comparison
Table 3 demonstrates the effectiveness of DAGCD across diverse QA datasets and model configurations. We summarize the findings based on datasets and model characteristics. Additionally, we evaluated DAGCD on summarization tasks, and the results demonstrate that DAGCD also achieves improvements (detailed results see Appendix F.6).
5.2.1 Dataset-Level Observations
DAGCD achieves consistent improvements across diverse QA tasks, including multi-hop reasoning, long-form retrieval, and document-level QA.
HotpotQA, TriviaQA, SearchQA: DAGCD demonstrates strong performance in tasks requiring multi-paragraph reasoning and long-form retrieval. It achieves the highest gains on HotpotQA with a 10.02% EM and 9.05% F1 improvement on Mistral-7B. Additionally, DAGCD outperforms CAD and COIECD on TriviaQA and SearchQA, with significant improvements across different models.
SQuAD, NewsQA, NQ: DAGCD demonstrates robust performance in single-paragraph and document-level tasks. On NQ, it achieves a 24.06% EM and 21.36% F1 increase on Mistral-7B compared to greedy decoding, while delivering consistent gains across SQuAD and NewsQA datasets.
NQ-Swap: In adversarial scenarios simulated by NQ-Swap, DAGCD achieves significant improvements, including 30.10% EM and 24.35% F1 gains on Mistral-7B, underscoring its robustness in challenging conditions.
5.2.2 Model-Level Observations
DAGCD demonstrates broad applicability across different model families, sizes, and tuning variants.
Model Families: DAGCD consistently improves performance across both LLaMA and Mistral families. On Mistral-7B, DAGCD achieves an average EM of 59.54 and an average F1 of 69.69, outperforming greedy decoding by 12.99% and 13.58%, respectively. LLaMA2 models also benefit significantly, on LLaMA2-7B, with average F1 improvements of 6.91% over greedy decoding.
Model Sizes: DAGCD improves performance across models of varying model sizes, achieving an average EM and F1 increase of 6.77% and 6.91% on LLaMA2-7B, and 5.63% and 6.35% on LLaMA2-13B, respectively.
Instruction-Tuned Models: DAGCD consistently achieves performance gains across instruction-tuned models, surpassing all baselines with the highest average EM and F1 scores.
Figure 4: The impact of different $\alpha$ values on the performance of our method (fixed top-rank = 10).
Figure 5: The impact of different Top-Rank constraint on the performance of DAGCD (fixed $\alpha=2$ and $4$).
5.3 Ablation Study
Impact of Different $\alpha$ We evaluated the impact of different values of $\alpha$ on model performance with the top-rank fixed at 10. The results, presented in Figure 4, indicate that $\alpha$ determines the adjustment intensity applied to the original generation distribution. For pretrained models, optimal performance is achieved at $\alpha=2$, whereas for Chat models, the best performance is observed at $\alpha=4$.
Impact of Different Top-Rank We evaluated F1 score variations under different top-rank filtering constraints on HotpotQA. As shown in Figure 5, top-rank filtering reduces false positives, with performance initially improving as constraints loosen, then declining when overly relaxed.
Additional results and performance variations under different prompts in Appendix F.7 and F.8.
6 Discussion and Analysis
6.1 Dynamic Decoding: Real-Time Efficiency Without Post-Generation Correction
Post-generation correction methods, such as CAD Shi et al. (2024b) and COIECD Yuan et al. (2024), improve contextual alignment but often rely on multi-step processes, resulting in significant computational overhead. In contrast, our dynamic decoding approach integrates context adjustments directly into the generation process, providing both efficiency and real-time optimization.
Lightweight Context Utilization Detector Using a logistic regression-based Context Utilization Detector, our method enables real-time adjustments with minimal computational overhead. This detector is efficient compared to resource-intensive alternatives, such as integrated gradients or exhaustive attention head manipulation.
Single-Pass Decoding with Real-Time Faithfulness Optimization By integrating the Context Utilization Detector directly into the decoding process, our method eliminates redundant steps such as output comparisons or external consistency checks. During generation, attention-based context utilization signals are leveraged in real time to proactively enhance faithfulness. This single-step strategy ensures that the output remains aligned with the input context without requiring additional post-processing, while maintaining the theoretical time complexity of regular decoding.
6.2 Interpretability Through Attention: Insights into Context Utilization
By systematically analyzing attention mechanisms, our approach uncovers how retrieved context influences the generation process and provides interpretable insights into the model’s behavior. Feature-Based Attention Analysis: Using naturally occurring cases (e.g., failure in closed-book settings but success in open-book settings), we systematically isolate attention patterns that are indicative of context utilization. A logistic regression model trained on these patterns identifies the most relevant attention heads with high accuracy, providing a quantitative measure of each head’s contribution to context alignment.
Transparent Decision-Making: The feature coefficients of the logistic regression model directly map to the importance of specific attention heads. This transparency allows for intuitive interpretation, making it clear which heads are most responsible for leveraging context tokens during generation.
7 Related Work
7.1 Context Faithfulness Hallucination
Current solutions to context faithfulness hallucination primarily focus on detection and mitigation. For detection, Lei et al. Lei et al. (2023) proposed a post-generation editing strategy that uses natural language inference to classify and revise hallucinated segments in generated content. Choi et al. Choi et al. (2023) introduced Knowledge-Constrained Decoding, which detects hallucinations at the token level during generation and reweights token distributions to guide output. Similarly, Chuang et al. Chuang et al. (2024a) proposed Lookback Lens Guided Decoding, which selects the most faithful output among multiple candidates to improve consistency.
For mitigation, CAD Shi et al. (2024b) compares outputs with and without concatenated context to enhance contextual adherence, while COIECD Yuan et al. (2024) improves upon CAD by incorporating entropy-based constraints to balance context usage. Wang et al. Wang et al. (2024) further introduced ADACAD, which dynamically adjusts token-level adherence using divergence between contextual and non-contextual outputs.
While effective, most existing methods face limitations such as high computational overhead and reliance on multiple decoding passes. In this work, we propose a real-time solution that integrates context utilization signals directly into the decoding process, achieving efficient and faithful generation without additional processing steps.
7.2 Attention Mechanisms and Interpretability
The attention mechanism provides valuable insights into how models prioritize different parts of an input sequence Clark et al. (2019a); Geva et al. (2023) and has become central to understanding Transformer-based LLMs Vashishth et al. (2019); Hao et al. (2021); Zhao et al. (2024). In LLMs, attention heads often perform distinct roles, such as capturing syntactic dependencies or aligning semantic relationships Olsson et al. (2022); Zheng et al. (2024); Jin et al. (2024b).
Recent studies have also explored the collaborative behavior of attention heads. For instance, the retrieval head framework Wu et al. (2024) identifies heads that collectively retrieve relevant tokens, while cutting-off-heads Jin et al. (2024b) highlights critical heads through systematic ablation. Gradient-based methods like IRCAN Shi et al. (2024a) further investigate the contributions of attention scores and neurons to model outputs.
Unlike prior work that focuses on individual heads or neurons, our study examines the collaborative patterns among multiple attention heads. By analyzing how attention mechanisms collectively utilize contextual tokens, we provide a holistic view of their role in aligning outputs with user-provided context, offering new insights into Transformer-based model behavior.
8 Conclusion
In this paper, we addresses context faithfulness hallucinations in LLMs by proposing Dynamic Attention-Guided Context Decoding (DAGCD), a lightweight framework that integrates attention distributions and entropy-based uncertainty signals to amplify contextually relevant tokens during generation. Our analysis revealed a strong correlation between high uncertainty and hallucinations, and probing experiments validated that attention mechanisms encode signals indicative of contextual utilization. Experiments across multiple QA datasets demonstrated that DAGCD achieves consistent improvements in context faithfulness, robustness, and scalability, providing an effective solution for context-sensitive generation tasks.
Limitations
Dependency on Classifier Accuracy and Robustness to Noisy Contexts
DAGCD relies on an attention-ratio based classifier to assess the relevance of context tokens during generation. While the classifier demonstrates high accuracy across datasets and models, its performance may degrade in scenarios with extremely long contexts, complex dialogues, or noisy inputs. Misclassifications in these cases could lead to incorrect adjustments, potentially amplifying irrelevant tokens or diminishing the contribution of critical ones. Similarly, the method’s robustness to adversarial or noisy contexts with misleading or irrelevant information remains an open challenge. Enhancing the classifier’s resilience and incorporating mechanisms to filter or downweight adversarial noise could further strengthen DAGCD’s applicability in real-world scenarios.
Scaling Factor Adjustment for Model Characteristics
The scaling factor $\alpha$ introduced in DAGCD needs to be adjusted based on the characteristics of different models. Although our study shows a strong correlation between entropy-guided uncertainty measures and the model’s uncertainty during generation, it does not establish a precise quantitative relationship. This limitation necessitates empirical calibration of the scaling factor for each model to ensure effective adjustments. Such calibration ensures that the method compensates for model-specific entropy variations, but it may introduce additional computational overhead during deployment.
Generalization Across Tasks and Domains
Our evaluation primarily focuses on QA tasks, leaving the generalization of DAGCD to other tasks, such as summarization or dialogue generation, unexplored. The attention-ratio based classifier, optimized for QA datasets, may require additional fine-tuning or redesign to handle different output structures and task-specific challenges. Extending the method to diverse domains and tasks could further validate its robustness and scalability.
Appendix A Prediction Uncertainty Analysis Details
A.1 Dataset and Sampling
We selected 6,000 samples from the MrQA training set Fisch et al. (2019), distributed evenly across six sub-datasets. Each sub-dataset contained 500 correctly answered samples and 500 incorrectly answered samples, evaluated using the Exact Match (EM) metric.
A.2 Prompt and Decoding Settings
The Prompt 1 template Yuan et al. (2024), as shown in Appendix G, was used for response generation. LLaMA2-7B Touvron et al. (2023) served as the backbone model, with greedy decoding ensuring deterministic outputs.
A.3 Computation Process
For each sample, we calculated two metrics to quantify uncertainty. First, the normalized entropy, defined in Equation 1, measures the dispersion of probabilities across the vocabulary, providing an overall view of the model’s uncertainty. Second, the top-1 token probability represents the likelihood of the most probable token, offering a complementary perspective on the model’s prediction confidence. These metrics focus on the initial generated token to analyze how uncertainty affects response faithfulness.
A.4 Dataset and Sampling
We selected 6,000 samples from the training set Fisch et al. (2019) for this analysis (The dataset consists of both the question and the context containing gold answer.). These samples were evenly distributed across six sub-datasets, with 1,000 samples drawn from each. For each sub-dataset, consisting of 500 correctly answered samples and 500 incorrectly answered samples (given context), the correctness of the responses is evaluated using the Exact Match (EM) metric.
A.5 Uncertainty Metrics Calculation Settings
We employed the Prompt 1 template Yuan et al. (2024) (as shown in Appendix G) for generation with LLaMA2-7B Touvron et al. (2023), using greedy decoding to ensure deterministic predictions. For each sample, we computed the normalized entropy of the distribution when the model generates the first answer token, while simultaneously recording the probability of the top-1 token in the distribution.
Appendix B Attention Signal Analysis Details
B.1 Dataset Construction
We constructed the dataset by selecting samples from the MrQA training set dataset Fisch et al. (2019), focusing on cases where the model’s output changed from incorrect to correct after context concatenation. A total of 2,000 positive samples (utilized tokens) and 2,000 negative samples (non-utilized tokens) were selected. And each sample consists of a label (utilized or non-utilized token) and the corresponding attention ratio feature vector $\mathbf{v}_{j}$ (under Prompt 1). The dataset was evenly split into training (50%) and testing (50%) sets.
B.2 Training and Evaluation
We trained a Logistic Regression (LR) classifier using 5-fold cross-validation. L2 regularization was applied to prevent overfitting. The classifier was evaluated across Transformer-based LLMs, including:
  • LLaMA2: 7B, 13B, 7B-Chat, 13B-Chat
  • Mistral: 7B, 7B-Instruct
Appendix C Details of LR Classifier Generalization and Feature Analysis
C.1 Cross-Prompt Test Results
To evaluate the robustness of the classifier under different prompts, we reconstructed the attention ratio feature vectors using Prompts 2, 3, and 4. These prompts differ in structure and phrasing but are consistent in task objectives. The classifier, trained using Prompt 1, was then tested on these alternative prompts.
The results, shown in Table 4, demonstrate that the classifier maintains an ACC exceeding 97% and AUC above 0.99 across all prompts. This indicates that the attention ratio signal is prompt-agnostic and generalizes well across different input structures.
Prompt LLaMA2-7B Mistral-7B
ACC AUC ACC AUC
Prompt2 0.9797 0.9932 0.9762 0.9902
Prompt3 0.9768 0.9926 0.9763 0.9889
Prompt4 0.9794 0.9946 0.9771 0.9927
Table 4: Performance testing under different prompts. Training data: attention ratio feature vector under Prompt 1. Test data: attention ratio feature vector under Prompts 2, 3, and 4.
Figure 6: Performance of the LR classifier with Top-$K$ and Bottom-$K$ features. Based on the absolute values of feature coefficients, the Top-$K$ and Bottom-$K$ features were selected to train an LR classifier with sparse features. The figure shows the ACC and AUC performance of the classifier on the test set for different values of $K$.
Figure 7: Out-of-domain performance of the LR classifier trained with full features and Top-10 features. Performance variations on out-of-domain data for LR classifiers trained using all features versus the top-10 features.
Figure 8: Heatmap of feature coefficients for LR. The heatmap of feature coefficients for LR classifiers trained using attention ratios from different LLMs as features.
C.2 Analysis of the Importance of Different Features
In analyzing the LR classifier trained with attention values from all attention heads as features, we found that most feature coefficients had small absolute values. This indicates that only a few attention heads are crucial for utilization detection.
To explore their impact, we selected the top-K and bottom-K features based on the absolute values of their coefficients and trained LR models using these subsets. Figure 6: subplots (a1) and (a2) shows how classification accuracy (ACC) changes with K. Using top-K features, the model achieves over 0.95 accuracy for all LLMs with K=10, matching the performance of using all features. In contrast, models with bottom-K features perform poorly, failing to reach 0.95 even with K=100. The AUC curves in Figure 6: subplots (b1) and (b2) for different K further confirm this. Models with top-K features maintain high accuracy and robustness, while those with bottom-K features show significantly worse performance. These results emphasize that a small number of key attention heads are enough for effective detection, while irrelevant features add little value and may introduce noise.
We also compared the performance of LR classifiers trained with all features versus only the top-10 features on out-of-domain data. As shown in Figure 7, the LR trained with only the top-10 features achieved better ACC and AUC on out-of-domain data.
Based on the ACC and AUC results, we find that the LR classifier trained with the Top-10 features achieves good accuracy and robustness while using the minimum number of features. Therefore, all subsequent inference employs LR classifiers trained with the Top-10 features.
Figure 9: Performance distribution of LR with single features. An LR classifier is trained using the attention ratio from a single head as features. The figure shows the ACC distribution of LR classifiers trained with attention ratios from different heads on the test set.
Figure 10: Performance of the LR classifier trained with Bottom-$K$ features. Based on the ACC of LR classifiers trained using the attention ratio from a single head as features, the Bottom-$K$ heads with the lowest ACC are selected. The LR classifier is then retrained using the attention ratios of these Bottom-$K$ heads as features, and its performance on the test set is presented.
Appendix D Analyzing Context Utilization Signal Contribution Across Heads
To further explore the role of attention heads in encoding contextual utilization signals, we conducted a detailed analysis to examine the strength and distribution of signals across heads. This section presents insights into the concentration of strong signals in certain heads, the independent utility of individual heads, and the complementarity of weaker heads.
D.1 Signal Strength and Concentration in Heads
To identify influential heads, we visualized the coefficients of the trained Logistic Regression (LR) classifier, which were derived from the attention ratio features of all heads. Approximately 5% of the heads exhibited significantly high coefficients, suggesting that these heads dominate the classification task (Figure 8). Repeating this analysis across 100 random seeds revealed consistent selection of these top heads, indicating their robustness as key signal carriers.
To evaluate the standalone utility of these heads, we trained LR classifiers using the attention ratio from a single head as the feature. The results in Figure 9 show that About 5% to 10% of the heads achieved classification accuracies (ACC) above 0.8, highlighting their ability to independently encode contextual utilization signals. However, the majority of heads performed poorly in isolation, with ACCs below 0.8. This disparity emphasizes the varying degrees of utility across heads, with a small subset contributing disproportionately strong signals.
Additionally, we also observed that on the Mistral model, the vast majority of heads perform well when acting individually. This indicates the presence of more high-performing heads in the Mistral model, which may explain why our method achieves greater improvements on Mistral compared to other models.
D.2 Complementary Contributions of Weaker Heads
Although most attention heads have limited standalone utility, we observe that combining weaker heads into subsets significantly improves classification performance. Figure 10 illustrates that we selected the bottom-K features, based on the classification accuracy of individual heads, to train the classifier and analyze the performance gain from combining weaker heads. The results show that when the number of bottom-K heads reaches 500, the classification accuracy stabilizes at approximately 90%. This finding highlights the complementarity of weaker heads, as their aggregated signals collectively achieve robust token classification.
D.3 Data-Dependent and Generalizable Heads
To investigate whether the influential heads identified in one dataset maintain their significance across other datasets, we conducted cross-dataset experiments. For datasets $A$, $B$, and $C$, we analyzed the top 10 heads ranked by LR coefficients and observed significant differences in head rankings across datasets. However, when considering the intersection of the top 100 heads across all three datasets, we identified a subset of heads that consistently exhibited high importance.
Using this subset of generalizable heads, we retrained the classifier on dataset $A$ and tested it on datasets $B$ and $C$. The model achieved high performance across all datasets, suggesting the presence of data-independent heads that encode universal contextual utilization signals. Conversely, dataset-specific top heads exhibited superior performance on their respective datasets, indicating a degree of data dependency in some heads’ contributions.
Summary of Findings
Our analysis reveals three key characteristics of attention heads in encoding contextual utilization signals:
  1. 1.
    Concentration: A small subset of heads consistently contributes strong independent signals, dominating the classification task.
  2. 2.
    Complementarity: Weaker heads collectively provide complementary signals, enabling robust classification when aggregated.
  3. 3.
    Data Dependency and Generalization: Some heads exhibit data-specific importance, while others demonstrate generalizability across datasets, reflecting a balance between dataset-specific and universal contributions.
These findings highlight the nuanced roles of attention heads in contextual token utilization and provide a foundation for further exploration of their properties and applications.
Appendix E Models and Datasets
E.1 Model Details
The LLMs used in this work, along with its HuggingFace ID, is as follows:
  • LLaMA2-7B: meta-llama/Llama-2-7b-hf
  • LLaMA2-7B-Chat: meta-llama/Llama-2-7b-chat-hf
  • LLaMA2-13B: meta-llama/Llama-2-13b-hf
  • LLaMA2-13B-Chat: meta-llama/Llama-2-13b-chat-hf
  • Mistral-7B: mistralai/Mistral-7B-v0.1
  • Mistral-7B-Instruct: mistralai/Mistral-7B-Instruct-v0.1
Appendix F Experimental Details
The datasets used in this study include seven open-book question-answering (QA) datasets, grouped into three categories based on their QA task characteristics: multi-hop reasoning, long-form retrieval-based QA, and single-paragraph extraction tasks. Additionally, an adversarial dataset is included for evaluating the robustness of the proposed method. Detailed descriptions and dataset statistics are provided below.
F.1 Dataset Categories and Statistics
Multi-Hop Reasoning (HotpotQA). HotpotQA Yang et al. (2018) is a benchmark dataset for multi-hop reasoning, requiring models to synthesize information across multiple paragraphs to generate an answer. This dataset emphasizes complex reasoning over distributed evidence, making it a critical benchmark for evaluating context utilization.
Long-Form Retrieval-Based QA (TriviaQA, SearchQA). TriviaQA Joshi et al. (2017) and SearchQA Dunn et al. (2017) require reasoning over longer contexts, with answers scattered across retrieved documents. These datasets test the model’s ability to focus on relevant content in lengthy contexts and generate precise answers.
Single-Paragraph Extraction (SQuAD). SQuAD Rajpurkar et al. (2016), and NewsQA Trischler et al. (2017) are standard extractive QA datasets where the answer is typically located within a single paragraph. These datasets are widely used for evaluating the span-extraction capabilities of QA systems.
Document-Level QA (NQ). NQ Kwiatkowski et al. (2019) is a document-level open-domain question answering dataset driven by real user queries. It requires systems to extract long answers from entire Wikipedia documents and generate specific short answers, evaluating document-level information retrieval and natural language understanding capabilities.
Simulated Conflict Scenarios (NQ-Swap). NQ-Swap Longpre et al. (2021) is an artificially constructed dataset that introduces adversarial entity swaps into NQ to create ambiguous or conflicting contexts. It evaluates the model’s ability to resolve conflicts and faithfully utilize context.
F.2 Dataset Sources and Formats
All datasets are standardized in the unified schema provided by the MrQA repository (Huggingface ID: mrqa-workshop/mrqa), except for NQ-Swap, which is sourced from a separate repository (Huggingface ID: pminervini/NQ-Swap). The datasets used for training the logistic regression model (§ 2) and attention analysis (§ 3) are drawn from the training sets of the MrQA repository. Model performance evaluation is conducted using the validation sets from the same repository. All datasets have been preprocessed to ensure compatibility with our experimental framework.
F.3 Dataset Statistics
Table 5 presents the size of the datasets used in this study to evaluate model performance.
Dataset Number of Samples
HotpotQA (Multi-Hop) 5904
TriviaQA (Long-Form Retrieval) 7785
SearchQA (Long-Form Retrieval) 16980
SQuAD (Single-Paragraph) 10507
NewsQA (Single-Paragraph) 4212
NQ (Document-Level) 12836
NQ-Swap (Simulated Conflicts) 4746
Table 5: Dataset statistics. A summary of the dataset sizes used for evaluation across different datasets.
F.4 Baseline Method Configurations
We implemented the baseline methods with their recommended hyperparameter settings for fair comparisons:
  • CAD Shi et al. (2024b): The contrastive adjustment factor $\alpha$ was set to 1.
  • COIECD Yuan et al. (2024): The entropy regularization parameter $\lambda$ was set to 0.25, and the contrastive adjustment factor $\alpha$ was set to 1.
F.5 Implementation Details
All experiments utilized a unified prompt template (Prompt 1) to ensure consistency across methods. The prompt format is detailed in Appendix G. For decoding, greedy decoding was employed to produce deterministic outputs and facilitate direct comparisons across methods. All models run on NVIDIA A100 GPUs.
Model Decoding ROUGE-L factKB BERT-P BERT-R BERT-F1
LLaMA2-7B Regular 0.2081 0.9932 0.9000 0.7997 0.8465
CAD 0.2361 0.9786 0.9054 0.8016 0.8514
COIECD 0.2089 0.9845 0.9152 0.8014 0.8543
OURs 0.2134 0.9856 0.9210 0.8026 0.8576
LLaMA2-7B-Chat Regular 0.2368 0.9846 0.9056 0.8035 0.8515
CAD 0.2082 0.9417 0.9001 0.7997 0.8466
COIECD 0.2371 0.9807 0.9055 0.8034 0.8513
OURs 0.2426 0.9866 0.9104 0.8036 0.8536
Table 6: Comparison of evaluation results on CNN/DailyMail. The table compares the evaluation results between greedy decoding and our proposed method on the CNN/DailyMail dataset. Bold denotes the best performance, while underlined indicates the second-best performance.
F.6 Results on Summarization Tasks
To validate the performance of our approach on long-form answer generation tasks, we conducted experimental evaluations on the CNN_DM See et al. (2017) summarization dataset (we randomly sampled 500 instances from the dataset for evaluation). Similar to prior work Shi et al. (2024b), we adopted ROUGE-L Lin (2004), factKB Feng et al. (2023), and BERTScore Zhang et al. (2020) as comprehensive evaluation metrics to assess both the accuracy and factual consistency of the generated content. The experimental results, as shown in Table 6, demonstrate that our method achieves significant improvements on both the pretrained and chat versions of LLaMA2.
F.7 Additional Results of The Impact of Different $\alpha$
We additionally evaluated the performance variations of our method on Mistral-7B and Mistral-7B-Instruct under different scaling factors $\alpha$. Figure 11 illustrates the performance changes on the HotpotQA dataset. For Mistral-7B, the optimal performance is achieved at $\alpha=5$. In contrast, for Mistral-7B-Instruct, the performance only stabilizes after $\alpha=13$. This indicates that different models may require different optimal scaling factors for the best performance.
F.8 Impact of Different Prompts
To assess robustness to prompt variations, we tested multiple prompts from prior studies Zhou et al. (2023); Yuan et al. (2024); Wang et al. (2024) (templates in Figure 13). Figure 12 illustrates the variations in F1 scores for LLaMA2-7B and Mistral-7B on the HotpotQA and NewsQA datasets under different prompt templates. The results show that DAGCD consistently outperforms baselines across all tested prompts, demonstrating its adaptability to diverse input formats and reliability across QA tasks.
Figure 11: The performance on HotpotQA for DAGCD under different scaling factors.
Figure 12: Performance variations across different prompt templates. The figure shows F1 score variations on the HotpotQA and NewsQA datasets for Greedy Decoding and OURs (DAGCD) under different prompt templates.
Appendix G Prompt Templates
Figure 13 shows the prompt templates used in this paper.
Figure 13: Prompt templates used in this paper.
Appendix H Symbol Explanation
The symbols and corresponding meanings defined in this paper, as shown in Table 7.
Symbol Meaning
$H_{\text{norm}}(P)$ Normalized entropy of distribution $P$
$v_{j}$ Attention ratio feature vector for token $j$
$v_{j}^{(K)}$ Top-K attention ratio feature vector for token $j$
$a_{l,h}^{j}$ The raw attention of the $h$-th head in the $l$-th layer for token $j$
$r_{l,h}^{j}$ The attention ratio of the $h$-th head in the $l$-th layer for token $j$
$s_{j}$ The utilization score for token $j$
$c_{k}$ The coefficient of the $k$-th feature in the LR classifier
$w_{k}$ The normalized coefficient of the $k$-th feature in the LR classifier
$U$ The utilization distribution
$u_{i}$ The normalized utilization score for token $j$
$U_{top}$ The subset of $U$ for tokens ranked in the top-$R$ of the generation distribution
Table 7: Annotation table.

### v2

@@
 Abstract
-Large language models (LLMs) often suffer from context faithfulness hallucinations, where outputs deviate from retrieved information due to insufficient context utilization and high output uncertainty.
-Our uncertainty evaluation experiments reveal a strong correlation between high uncertainty and hallucinations.
-We hypothesize that attention mechanisms encode signals indicative of contextual utilization, validated through probing analysis.
-Based on these insights, we propose Dynamic Attention-Guided Context Decoding (DAGCD), a lightweight framework that integrates attention distributions and uncertainty signals in a single-pass decoding process.
-Experiments across QA datasets demonstrate DAGCD’s effectiveness, achieving significant improvements in faithfulness and robustness while maintaining computational efficiency.
+Large language models (LLMs) often exhibit Context Faithfulness Hallucinations, where outputs deviate from retrieved information due to incomplete context integration.
+Our analysis reveals a strong correlation between token-level uncertainty and hallucinations.
+We hypothesize that attention mechanisms inherently encode context utilization signals, supported by probing analysis.
+Based on these insights, we propose Dynamic Attention-Guided Context Decoding (DAGCD), a lightweight framework that leverages attention distributions and uncertainty signals in a single-pass decoding.
+Experiments on open-book QA datasets demonstrate DAGCD’s effectiveness, yielding significant improvements in faithfulness and robustness while preserving computational efficiency.
+[^1] Our code is available at uestc-huangyw/DAGCD.
 Dynamic Attention-Guided Context Decoding for Mitigating Context Faithfulness Hallucinations in Large Language Models
@@
 (2023); Touvron et al.
-(2023) excel at generating fluent and contextually relevant responses.
-However, they often struggle with producing factually accurate outputs, particularly when relying on external information Vu et al.
+(2023) excel in generating fluent and contextually relevant responses.
+However, they often struggle with factual accuracy, especially when relying on external information.
+Vu et al.
 (2023).
@@
 (2023).
-These issues undermine RAG systems’ reliability, particularly in high-stakes applications requiring factual accuracy Chuang et al.
+These hallucinations undermine the reliability of RAG systems, particularly in critical domains where factual accuracy is paramountChuang et al.
 (2024a).
-Figure 1: Analysis of generation distribution when concatenating context: ($a_{1},a_{2}$) the model’s uncertainty when generating correct versus wrong answers; ($b$) for wrong answers, the ranking of the gold answer token in the distribution, and the probability difference between the gold answer token and the Top-1 predicted token.
 Existing methods, such as CAD Shi et al.
 (2024b) and COIECD Yuan et al.
-(2024), attempt to mitigate context faithfulness hallucinations by dynamically adjusting decoding distributions through generation distribution comparisons or token-level uncertainty signals.
-While effective to some extent, these methods face several key limitations: limited interpretability, degraded performance when context-agnostic and context-aware outputs differ significantly, and high computational overhead due to multiple decoding passes.
-These issues highlight the need for lightweight and interpretable approaches that can effectively utilize retrieved context while maintaining scalability.
-To further explore these issues, we analyze the relationship between output uncertainty and context faithfulness hallucinations.
-As shown in Figure 1, we observe a strong correlation between context hallucinations and high entropy in the generation distribution.
-Specifically, incorrect generations often reflect significant uncertainty through high entropy.
-Notably, in cases of wrong answers, context tokens corresponding to the correct answer often appear within the top-10 ranks but are assigned insufficient confidence.
-This indicates that the model’s internal mechanisms encode interpretable signals to indicate context utilization.
-Building on our analysis and findings in interpretability studies, we hypothesize that attention mechanisms in Transformer models encode signals indicative of contextual utilization.
-Attention modules play a critical role in aligning generated tokens with their corresponding context tokens by distributing focus across input tokens Geva et al.
-(2023); Jin et al.
-(2024b).
-To validate our hypothesis, we trained a probing classifier using Logistic Regression on attention distributions, achieving over 97% accuracy in distinguishing contextually relevant tokens.
-These results demonstrate that attention weights can serve as a lightweight and interpretable proxy for assessing the effective incorporation of retrieved evidence during generation.
-Motivated by these findings, we propose Dynamic Attention-Guided Context Decoding (DAGCD), a novel algorithm to mitigate context faithfulness hallucinations.
+(2024), attempt to mitigate context faithfulness hallucinations by dynamically adjusting decoding distributions through token-level probability distribution comparisons or token-level uncertainty signals.
+While effective to some extent, these methods face several key limitations: limited interpretability, degraded performance when context-agnostic and context-aware outputs differ significantly, and computational complexity due to multiple decoding passes.
+Figure 1: Analysis of the token-level probability distribution after context concatenation: (a1, a2) the model’s uncertainty when generating correct versus wrong answers, measured by NE and MSP; (b) for wrong answers, the ranking of the golden answer token within the token-level probability distribution.
+To better understand context faithfulness hallucinations and explore potential solutions, we take an internally-driven approach, analyzing intrinsic model signals that may explain why retrieval-augmented large language models struggle to utilize retrieved context faithfully Liang et al.
+(2024).
+Motivated by prior research linking token-level uncertainty to factual hallucinations Chuang et al.
+(2024b); Das et al.
+(2025), we examine whether token-level probability distribution entropy correlates with context faithfulness hallucinations.
+Our analysis reveals a strong correlation between higher uncertainty and context faithfulness hallucinations.
+Specifically, as shown in Figure 1 (a1), wrong answers exhibit higher entropy in the token-level probability distribution, indicating greater uncertainty in generation.
+Even for correct answers (Figure 1 (a2)), it often assigns low confidence to the highest-ranked token, suggesting incomplete integration of retrieved context.
+Notably, in incorrect responses, most gold answer tokens appear in the top 10 in the token-level probability distribution, but are not assigned the highest probability (Figure 1 (b)), implying that the model identifies relevant context but fails to prioritize it effectively.
+These findings indicate that while models retrieve relevant context, they struggle to integrate and prioritize it during generation.
+Since most gold answer tokens appear within the top 10, an effective strategy to mitigate context faithfulness hallucinations is to dynamically identify and prioritize these tokens during generation.
+This requires detecting reliable signals that indicate how retrieved context influences the model’s predictions.
+Attention mechanisms in Transformer models naturally emerge as a key source of such signals, since they facilitate information flow between tokens Olsson et al.
+(2022a); Meng et al.
+(2022a); Geva et al.
+(2023).
+We hypothesize that attention distributions encode intrinsic indicators of context utilization.
+To validate our hypothesis, we trained a probing classifier using Logistic Regression on attention distributions, achieving over 0.99 AUC in distinguishing contextually relevant tokens.
+Even with just 100 training samples, the classifier demonstrated strong generalization across in-domain and cross-domain test sets, indicating that attention distributions inherently encode context utilization signals.
+These results support attention-based context utilization as a fundamental mechanism in LLMs.
+By leveraging these intrinsic signals, attention distributions provide a lightweight and interpretable means to assess how models integrate retrieved context into their predictions.
+Motivated by these findings, we propose Dynamic Attention-Guided Context Decoding (DAGCD), a novel method to mitigate context faithfulness hallucinations.
 Inspired by the copy-generator framework See et al.
 (2017); Xu et al.
-(2020b), DAGCD integrates attention weights to estimate the relevance of tokens in the retrieved context, dynamically adjusting output probabilities.
-Entropy-based uncertainty measures further refine these adjustments, amplifying underconfident but contextually relevant tokens.
-By unifying these strategies, DAGCD aligns outputs with context while maintaining computational efficiency.
+(2020), DAGCD integrates attention weights to estimate the relevance of tokens in the retrieved context, dynamically adjusting output probabilities.
+Additionally, token-level uncertainty guides these adjustments by emphasizing underconfident yet contextually relevant tokens.
+By combining these strategies, DAGCD ensures output alignment with the retrieval context and maintains efficiency.
 Our contributions are as follows:
   1. 1.
-    Comprehensive analysis of context faithfulness hallucinations: We reveal a strong correlation between context faithfulness hallucinations and output uncertainty through entropy-based analysis, revealing the model’s challenges in aligning with retrieved context under high uncertainty.
+    Comprehensive analysis of context faithfulness hallucinations: We identify a strong correlation between token-level uncertainty and context faithfulness hallucinations, showing that incorrect responses exhibit higher entropy and retrieved context is often recognized but not effectively prioritized.
   2. 2.
@@
   3. 3.
-    Lightweight and efficient decoding: DAGCD operates in a single decoding pass, integrating attention signals and uncertainty measures without additional overhead, achieving improved decoding efficiency.
+    Lightweight and efficient decoding: DAGCD operates in a single decoding pass, integrating attention signals and uncertainty measures without additional overhead, improving efficiency.
   4. 4.
-    Extensive validation across datasets and models: DAGCD outperforms greedy decoding across multiple QA datasets, achieving an average F1 score improvement of 8.95% on pretrained models and 1.91% on instruction-tuned models, demonstrating its robustness and scalability.
+    Extensive validation across datasets and models: DAGCD outperforms greedy decoding across multiple QA datasets, improving EM by 17.67% on pretrained models and 2.25% on instruction-tuned models, demonstrating robustness and scalability.
 2 Why Can’t Generate Faithful Answers?
-LLMs often fail to generate responses that are faithful to the concatenated context.
-We analyzed (1) the impact of context concatenation on output uncertainty (§2.1) and (2) the ranking of contextual answer tokens within predictions (§2.2).
-Our findings indicate that high output uncertainty and low confidence in contextually relevant tokens are critical factors contributing to this issue.
+Token-level uncertainty is closely related to factual hallucinations, as models often exhibit higher entropy in the token-level probability distribution when generating factually incorrect outputs Chuang et al.
+(2024b); Das et al.
+(2025).
+While uncertainty measures help detect hallucination-prone predictions, most studies focus on factual hallucinations, where responses are incorrect without retrieved context.
+In contrast, context faithfulness hallucinations arise when models rely on retrieved information but generate misaligned or contradictory outputs.
+Despite their significance in RAG, their relationship with uncertainty remains unclear.
+Inspired by prior research, we investigate whether unfaithful responses in RAG exhibit higher entropy and whether contextually relevant tokens are recognized but assigned insufficient confidence.
 2.1 Uncertainty Leads to Unfaithful Answers
-To assess the relationship between uncertainty and response faithfulness, we focus on two widely used metrics: entropy, which measures the overall uncertainty in the output distribution Xu et al.
-(2020a); van der Poel et al.
-(2022); Huang et al.
-(2023b), and top-1 token probability, which reflects the model’s confidence in its highest-ranked prediction Yao et al.
-(2023); Luo et al.
-(2024).
-Higher entropy indicates greater uncertainty, while higher top-1 probability corresponds to greater confidence in the model’s predictions.
-Experimental Setup We compute the normalized entropy and top-1 token probability for the first generated token. Normalized entropy is defined as:
-$H_{\text{norm}}(P)=-\frac{\sum_{i=1}^{N}P_{i}\log P_{i}}{\log N},$ (1)
-where $P$ represents the probability distribution, and $N$ denotes the vocabulary size.
-These metrics quantify prediction confidence, with lower entropy and higher top-1 probability reflecting greater certainty.
-Detailed descriptions of the datasets, models, and calculation processes are provided in Appendix A.
-Results and Analysis Figure 1 (subplots $a_{1}$ and $a_{2}$) illustrates the normalized entropy and top-1 token probability distributions for correct and wrong answers.
-The model exhibits higher uncertainty for wrong answers, with an average normalized entropy of 0.36 compared to 0.29 for correct cases, and a lower average top-1 probability of 0.25 compared to 0.41.
-Notably, a significant overlap exists between the entropy distributions of correct (blue) and wrong (red) cases, indicating that even correct answers often exhibit high uncertainty.
-Model w/o context w/ context
-LLaMA2-7B -0.43 -0.53
-LLaMA2-7B-Chat -0.27 -0.33
-LLaMA2-13B -0.30 -0.51
-LLaMA2-13B-Chat -0.26 -0.32
-Mistral-7B -0.23 -0.22
-Mistral-7B-Instruct -0.30 -0.33
-Table 1: Spearman correlation analysis.
-We examine the relationship between F1 scores (pred-ans vs.
-gold-ans) and norm-entropy (generation distribution) under both w/o and w/ context settings ($p\ll 0.05$ for all models).
-Building on this observation, we conduct a Spearman correlation analysis (Table 1), further demonstrating how uncertainty (norm-entropy of the generation distribution) negatively correlates with accuracy (F1 score between predicted and golden answers) across all models.
-After concatenating context, this correlation becomes even more pronounced, confirming that uncertainty is a key factor in generating unfaithful answers.
+Experimental Setup
+To assess the relationship between uncertainty and response faithfulness, we use two common metrics: Normalized Entropy (NE), which measures the overall uncertainty in the token-level probability distribution Huang et al.
+(2023b), and Maximum Softmax Probability (MSP), which quantifies the model’s confidence in the highest probability within the token-level probability distribution Hendrycks and Gimpel (2017).
+Higher NE indicates greater uncertainty, while higher MSP corresponds to greater confidence in the predictions of the model.
+For detailed experimental descriptions, see Appendix A.
+Results and Analysis
+Figure 1 (a1) and (a2) illustrates the Normalized Entropy and Maximum Softmax Probability of the token-level probability distribution when the model produces correct and wrong answers.
+The model exhibits higher uncertainty for wrong answers, with an average NE of 0.36 compared to 0.29 for correct cases, and a lower average MSP of 0.25 compared to 0.41.
+Notably, there is a substantial overlap between the correct (blue) and wrong (red) cases in the figure, indicating that even correct answers often exhibit high uncertainty.
+We also analyzed the correlation between prediction accuracy and uncertainty, and the results demonstrate a significant negative correlation, further confirming that token-level uncertainty is strongly associated with unfaithful answers.
+For detailed results see Appendix A.4.
 2.2 LLM is Actually Utilizing Context
-Experimental Setup Using the wrong answer samples from §2.1, we analyze the ranking distribution of the gold answer token in the first generated distribution following context concatenation.
-Ranks are grouped into intervals (e.g., rank 2–4, rank 5–7, etc.), and we compute the proportion of gold answer tokens within each rank interval.
-Additionally, we calculate the average probability gap between the gold answer token and the top-1 token.
-Results and Analysis As shown in Figure 1 (subplot $b$), when the model answers incorrectly after context concatenation, 43% of the samples have the gold answer token ranked between 2–4, and a total of 66% have the gold answer token ranked within the top 10.
-Moreover, the average probability gap between the gold answer token and the top-1 token is relatively small: only 0.14 for ranks between 2–4 and 0.24 for ranks greater than 30.
-This indicates that, in most cases, the model can correctly identify the gold answer token within the context after concatenation, but it often fails to prioritize it effectively during the generation process.
-The findings highlight that in context faithfulness hallucination scenarios, the model utilizes correct context tokens but assigns them insufficient confidence, limiting their influence on the output.
-This limitation underscores the model’s incomplete integration of retrieved context, prompting an exploration of how context tokens are incorporated to guide strategies for enhancing faithfulness and mitigating hallucinations.
+Our previous analysis links token-level uncertainty to unfaithful responses, showing that incorrect outputs often have higher entropy and lower confidence.
+However, this does not mean the model ignores retrieved context.
+A key question remains: Do incorrect responses imply that LLMs have failed to leverage the retrieved context?
+To investigate this, we examine the ranking of gold answer tokens in the token-level probability distribution for incorrect responses.
+If these tokens frequently rank high but are not assigned the highest probability, it suggests the model identifies relevant context but fails to prioritize it effectively.
+Experimental Setup
+We analyze incorrect responses from §2.1 by examining the ranking distribution of gold answer tokens in the token-level probability distribution of the first generated token after context concatenation.
+The ranks are grouped into intervals, and we compute the proportion of gold answer tokens within each rank interval.
+Results and Analysis
+As shown in Figure 1 (subplot b), when the model generates an incorrect response, 43% of cases have the gold answer token ranked 2nd to 4th based on token probabilities, and 66% have it ranked within the top 10 in the token-level probability distribution.
+Moreover, as shown in Figure 6, the average probability gap between the gold answer token and the highest-probability token remains relatively small: 0.14 for ranks between 2 and 4, and 0.24 for ranks beyond 30.
+These findings indicate that in context faithfulness hallucination scenarios, the model recognizes relevant context tokens but fails to prioritize them effectively, limiting their impact on the generated output.
+This highlights an incomplete integration of retrieved context, emphasizing the need for improved context incorporation strategies to enhance faithfulness and mitigate hallucinations.
 3 Context Utilization Signal in Attention
-3.1 Background and Hypothesis
-Building on our findings, we focus on the attention mechanisms in Transformer models as a potential indicator of context token utilization.
-Attention mechanisms distribute focus across input tokens during generation, capturing the alignment between generated tokens and their corresponding context tokens Clark et al.
+Our analysis in Section 2 highlights that while models often identify relevant context tokens, they fail to assign them sufficient confidence, leading to context faithfulness hallucinations.
+To better understand this phenomenon, we seek reliable signals that indicate which retrieved context tokens are effectively utilized by the model during generation.
+Due to their role in integrating and propagating information across tokens, attention mechanisms naturally emerge as a key candidate for capturing such signals Olsson et al.
+(2022a); Meng et al.
 (2022a); Geva et al.
 (2023).
-We hypothesize that attention heads encode signals indicating the contribution of retrieved context tokens to the generation of specific outputs.
-This provides a lightweight and interpretable approach to understanding contextual integration.
+We hypothesize that attention distributions encode intrinsic indicators of context utilization, providing a lightweight and interpretable means to assess how models incorporate retrieved context into their outputs.
 3.1 Attention Ratio
-Analyzing raw attention weights poses challenges, as attention sink often results in disproportionate focus on non-context tokens such as delimiters, introducing noise to the analysis.Xiao et al.
-(2024); Clark et al.
-(2019b); Bondarenko et al.
-(2021); Son et al.
-(2024); Sun et al.
+A key challenge in analyzing attention weights is the noise from non-context tokens (e.g., delimiters) caused by attention sink effects Bondarenko et al.
+(2021); Xiao et al.
 (2024).
-Additionally, varying attention magnitudes across heads and layers Jain and Wallace (2019); Vig and Belinkov (2019) complicate feature comparison.
-To address these challenges, we compute normalized attention ratios exclusively for context tokens, reducing noise and stabilizing values across layers and heads.
-This approach constrains features to $[0,1]$, enabling consistent evaluation of context token importance.
-Formally, the attention ratio for token $j$ during the generation step, at the $h$-th head in the $l$-th layer, is defined as:
+Additionally, attention magnitudes vary across heads and layers, complicating feature comparison Jain and Wallace (2019); Vig and Belinkov (2019).
+To address these issues, we introduce the attention ratio, a normalized measure that captures how much attention a retrieved context token receives relative to the total attention assigned within the context.
+For a given token $j$ in the retrieved context $C$, the attention ratio at the $h$-th head in the $l$-th layer is defined as:
 $r_{l,h}^{j}=\frac{a_{l,h}^{j}}{\sum_{j\in C}a_{l,h}^{j}}$ (1)
-where $a_{l,h}^{j}$ represents the raw attention weight assigned to token $j$ in the context $C$, and $\sum_{j\in C}a_{l,h}^{j}$ is the total attention weight for all tokens in $C$.
-This ratio quantifies the relative importance of token $j$ within the context for a specific attention head.
-For each token $j$, we construct a feature vector by aggregating its attention ratios across all heads:
+where $a_{l,h}^{j}$ denotes the raw attention weight assigned to token $j$. This ratio quantifies the relative importance of $j$ within the retrieved context for a specific attention head.
+To construct token-level features, we aggregate attention ratios across all heads:
 $v_{j}=\left[r_{1,1}^{j},\dots,r_{\text{num\_layers},\text{num\_heads}}^{j}\right]$ (2)
-the attention ratio feature vector $v_{j}$ provides a comprehensive representation of the token’s attention patterns, capturing its relative importance across the model’s entire architecture.
-3.3 Experimental Setup and Results
-To validate our hypothesis, we focused on cases where the model’s output shifted from incorrect to correct after context concatenation.
+This feature vector represents the distribution of attention across layers and heads, enabling a structured assessment of context token importance.
+Figure 2: Cross Domain Validation. One dataset as the training set (X-axis) and the remaining datasets as the test sets, showing the AUC on the test sets (Y-axis).
+3.2 Experimental Setup
+To validate our hypothesis and investigate whether this mechanism generalizes across different datasets, training sizes, and prompt templates, we conduct the following experiments.
+Data Construction We constructed the dataset by randomly selecting samples from the MrQA training set dataset (contains six open-book QA datasets in different domains) Fisch et al.
+(2019), focusing on cases where the model’s output changed from incorrect to correct after context concatenation Meng et al.
+(2022b).
+These cases indicate that the model successfully leveraged the retrieved context to produce the correct answer.
 Context tokens were labeled as positive (utilized) if they corresponded to the gold answer, and negative (non-utilized) otherwise.
-Using these labels, we extracted attention ratio feature vectors $\mathbf{v}_{j}$ under Prompt 1 to train a Logistic Regression (LR) classifier.
-The classifier was evaluated across multiple Transformer-based LLMs, including LLaMA2 and Mistral models.
-Details on dataset construction, training, and evaluation are in Appendix B.
-Model ACC AUC
-LLaMA2-7B 0.9793 0.9925
-LLaMA2-7B-Chat 0.9798 0.9970
-LLaMA2-13B 0.9740 0.9935
-LLaMA2-13B-Chat 0.9737 0.9924
-Mistral-7B 0.9787 0.9915
-Mistral-7B-Instruct 0.9756 0.9883
-Table 2: Performance of the LR classifier. We report the average performance across different random seeds.
-The LR classifier demonstrates excellent performance across all evaluated LLMs, achieving accuracy above 0.97 and AUC scores exceeding 0.99 (Table 2).
-These results confirm that attention ratio features effectively distinguish utilized tokens from non-utilized ones, validating the hypothesis that attention patterns encode signals of contextual utilization.
-This robust performance, observed across various LLaMA2 and Mistral models, including chat and instruction-tuned variants, highlights the generality of attention ratio features across different architectures and fine-tuning paradigms.
-Moreover, the findings emphasize the potential of attention-based features as lightweight, interpretable signals for improving RAG systems.
-3.4 Generalization and Feature Analysis
-To validate the robustness and generalization of the LR classifier, we conducted extensive evaluations across out-of-domain data, prompt variations, and feature contributions.
-Out-of-Domain Evaluation We tested the classifier on six sub-datasets from different domains, training on five sub-datasets and evaluating on the held-out one.
-As shown in Figure 2, the classifier achieves an average accuracy (ACC) and area under the curve (AUC) above 0.95 across all datasets and LLMs.
-These results demonstrate strong generalization capabilities of attention ratio features, even when applied to unseen domains.
-Additional Analyses We further evaluated the robustness of attention-based features to prompt variations and analyzed the importance and characteristics of individual features.
-Results show that the classifier consistently performs well across prompts with minimal variation and the performance of classifier with top-K features surpasses that of using full features.
-Additionally, attention heads exhibit complementary behavior and data-dependent characteristics.
-Detailed results for these analyses are provided in Appendix C and Appendix D.
-Figure 2: Out-of-domain performance of the LR classifier. One dataset is used as the test set, and the remaining datasets are used as the training set.
+Using these labels, we extracted attention ratio feature vectors $\mathbf{v}_{j}$ to train a Logistic Regression (LR) classifier.
+3.3 Results
+Cross Domain Validation
+We tested the classifier on six sub-datasets from different domains.
+Specifically, we selected one dataset as the training set and tested the performance on the remaining datasets (each sub-dataset construct 500 samples, contains 250 positive and 250 negative samples).
+As shown in Figure 2, the classifier achieves an average AUC above 0.99 across all datasets and LLMs. The results indicate that the context utilization signal in attention is data-independent.
+Training Data Size Validation
+Building on the cross domain experiment, we further tested the impact of training set size on classifier performance.
+Specifically, we trained the model using data constructed from one sub-dataset and evaluated its performance on the remaining sub-datasets.
+Figure 3 shows the variation in model AUC with different training data sizes (where "Train size = 100" refers to a training set constructed with 50 positive and 50 negative samples).
+The results indicate that even with only 100 samples, the model’s AUC exceeds 0.96, and the performance improvement becomes limited as the data size increases.
+This demonstrates that the LR classifier trained using the attention ratio exhibits strong data-efficiency.
+Figure 3: Training Data Size Validation. Training sets of varying sizes were constructed from a single dataset (HotpotQA), and evaluate on remain datasets.
+Additional Results
+We then examined the impact of different prompt templates, and the results indicate that the classifier consistently maintains high performance regardless of the prompt template used.
+For Details, see Appendix B.
+To better apply the classifier to practical tasks, we conducted a detailed analysis of the importance of different features and the relationships between them.
+The results show that the classifier using the top-K features outperforms the one using the full feature set.
+Furthermore, the attention heads exhibit Concentration and Complementarity features.
+For Details, see Appendix C.
+Conclusion: A Fundamental Mechanism in Transformer-based LLMs
+The consistent generalization of attention-based context utilization across datasets, data sizes, and prompts reinforces its role as a fundamental mechanism in LLMs.
+Our findings show that attention heads encode robust context integration signals, providing a lightweight and interpretable way to assess how models incorporate retrieved context.
 Figure 4: The illustration of the generation process of our proposed DAGCD method.
@@
 (2017); Xu et al.
-(2020b), we propose Dynamic Attention-Guided Context Decoding (DAGCD) to migrate context faithfulness hallucinations.
+(2020), we propose Dynamic Attention-Guided Context Decoding (DAGCD) to mitigate context faithfulness hallucinations.
 DAGCD leverages utilization signals to dynamically guide the generation process, focusing on relevant contextual tokens.
-It integrates three steps: detecting utilized context tokens during inference (§4.1), constructing a utilization distribution (§4.2), and adjusting the generation probabilities to enhance contextual relevance (§4.3).
+It integrates three steps: detecting utilized context tokens during inference (§4.1), constructing a utilization distribution (§4.2), and adjusting the token-level probability distribution to enhance contextual utilization (§4.3).
 4.1 Context Utilization Detection
-Context Utilization Detector In Section 3.4, we observed that the LR classifier trained using attention ratios from all attention heads did not generalize as effectively as the LR classifier trained with top-K features.
-Consequently, we selected the LR model trained with top-K features as the Context Utilization Detector.
-The set of attention heads corresponding to the top-K features used for training is denoted as $A=[A_{l_{1},h_{1}},A_{l_{2},h_{2}},\dots,A_{l_{K},h_{K}}]$, where each head $A_{l_{k},h_{k}}$ is associated with a feature coefficient $c_{k}$ (for all models, we set $K$ = 10).
-[…]
-$\mathbf{v}_{j}^{(K)}=[r_{l_{1},h_{1}}^{j},r_{l_{2},h_{2}}^{j},\dots,r_{l_{K},h_{K}}^{j}]$ (4)
+Context Utilization Detector To identify contextually relevant tokens during inference, we utilize a Logistic Regression (LR) classifier trained on attention-based utilization signals.
+Our analysis in Section 3.3 shows that selecting the most informative attention heads improves generalization.
+Thus, we construct the Context Utilization Detector based on the top-K most important attention heads.
+Feature Data Collection To obtain feature vectors, we extract attention distributions at the current decoding step, as illustrated in Figure 4.
+Specifically, we take the last row of the attention map for each selected head $h_{k}\in H$ ($H$ is the set of top-K attention heads).
+To ensure focus on relevant information, we filter out non-contextual tokens, such as query tokens and placeholder tokens in templates.
+The top-K feature vector for each contextual token $j$ is then constructed as:
+$\mathbf{v}_{j}^{(K)}=[r_{h_{1}}^{j},r_{h_{2}}^{j},\dots,r_{h_{K}}^{j}]$ (3)
+where $r_{h_{k}}^{j}$ represents the normalized attention ratio of token $j$ in attention head $h_{k}$.
 Finally, the feature vector $\mathbf{v}_{j}^{(K)}$ is fed into the detector, which identifies the set of context tokens actively utilized at the current decoding step.
 4.2 Utilization Distribution Construction
-The context utilization detector identifies utilized tokens but does not quantify the degree of utilization for each token.
-To address this, we aggregate the attention ratios from different attention heads, weighted by their normalized feature coefficients $w_{k}$, to compute a utilization score for each token.
-Formally, the utilization score $s_{j}$ for token $j$ is defined as:
-$s_{j}=\sum_{k=1}^{K}(r_{l_{k},h_{k}}^{j}\times w_{k}),\ w_{k}=\frac{c_{k}}{\sum_{k=1}^{K}c_{k}}$ (5)
-The feature coefficient reflects the significance of each attention head, with positive values indicating a higher likelihood of token utilization and negative values suggesting a reduced likelihood.
-The utilization distribution is then constructed based on the utilization scores of individual tokens:
+The context utilization detector identifies which tokens are utilized but does not quantify the degree of utilization for each token.
+To address this, we compute a utilization score $s_{j}$ for each token $j$ by aggregating attention ratios from selected attention heads, weighted by their normalized feature coefficients $w_{k}$.
+Tokens classified as unused by the detector are directly assigned a score of zero.
+$s_{j}=\sum_{k=1}^{K}(r_{h_{k}}^{j}\times w_{k}),\ w_{k}=\frac{c_{k}}{\sum_{k=1}^{K}c_{k}}$ (4)
+where $r_{h_{k}}^{j}$ is the normalized attention ratio of token $j$ in attention head $h_{k}$, and $w_{k}$ is the importance weight assigned to head $h_{k}$.
+The coefficient $c_{k}$ is learned from the LR classifier, representing the contribution of each head to context utilization.
+The utilization distribution $U$ represents a probability distribution over context tokens, normalized based on their utilization scores:
 $\mathbf{U}=[u_{1},u_{2},\dots,u_{N}],\quad u_{i}=\frac{s_{i}}{\sum_{j=1}^{N}s_{j}}$ (5)
-where $u_{i}$ represents the normalized utilization score for token $i$.
-Here, $N$ denotes the size of the vocabulary, and $s_{i}=0$ for tokens either absent from the context or not utilized, resulting in $u_{i}=0$ for such tokens.
-4.3 Generate More Faithful Answers
-Building on the utilization signals and the utilization distribution, DAGCD refines the generation process by dynamically adjusting token probabilities to prioritize contextually relevant tokens.
-[…]
-Dynamic Scaling of Adjustments The intensity of adjustments is guided by the model’s output uncertainty, measured as the normalized entropy $H_{\text{norm}}(P)$ of the generation distribution, as defined in Equation 1.
+where $u_{i}$ denotes the utilization probability of token $i$, $N$ is the vocabulary size.
+Tokens either absent from the context or classified as non-utilized by the detector ($s_{i}=0$) are assigned $u_{i}=0$.
+Top-Rank Constraint
+To enhance the reliability of generation adjustments, our approach applies a top-rank restriction, ensuring modifications focus on plausible tokens.
+Specifically, we define $U_{\text{top}}$ as the subset of the utilization distribution $U$ corresponding to tokens ranked within the top-$R$ positions of the generation distribution.
+This design builds on prior work Li et al.
+(2023); Chuang et al.
+(2024b), addressing context faithfulness hallucination challenges while leveraging our observation that correct context tokens usually appear within the top-ranked positions in the token-level probability distribution, even when the model generates incorrect answers.
+Through this constraint, we reduce the risk of amplifying irrelevant or nonsensical tokens, preserving output integrity.
+4.3 Generating More Faithful Answers
+DAGCD adjusts token probabilities based on token-level uncertainty, measured using the normalized entropy $H_{\text{norm}}(P)$ of the token-level probability distribution.
 High entropy indicates greater uncertainty and a higher risk of generating contextually inconsistent responses.
-Since our analysis shows that entropy correlates with uncertainty but does not establish a direct numerical relationship, we introduce a scaling factor $\alpha$, which compensates for model-specific entropy variations.
-Top-Rank Filtering and Distribution Adjustment To enhance the reliability of generation adjustments, our approach incorporates a top-rank restriction, ensuring modifications focus on plausible tokens.
-Specifically, we define $U_{\text{top}}$ as the subset of the utilization distribution $U$ corresponding to tokens ranked within the top-$R$ positions of the generation distribution.
-This design addresses challenges similar to those highlighted in prior work Li et al.
-(2023); Chuang et al.
-(2024b), while specifically leveraging the observation that correct context tokens are typically found within the top-rank set, even when the model produces incorrect answers.
-Restricting adjustments to $U_{\text{top}}$ reduces the risk of amplifying irrelevant or nonsensical tokens, thereby maintaining the output’s integrity.
-Under this framework, adjustments are applied only when utilized tokens in $U_{\text{top}}$ overlap with the top-rank set of the generation distribution.
+Since entropy correlates with uncertainty but lacks a fixed numerical relationship, we introduce a scaling factor $\alpha$ to compensate for model-specific entropy variations.
+Adjustments are applied only when utilized tokens in $U_{\text{top}}$ overlap with the top-ranked tokens in the token-level probability distribution.
 The adjusted generation distribution $P^{\prime}$ is computed as:
 $P^{\prime}=P+\alpha H_{\text{norm}}(P)\cdot U_{\text{top}}$ (6)
-where $P$ is the original generation distribution, and $\alpha H_{\text{norm}}(P)$ scales the adjustment intensity dynamically based on the normalized entropy $H_{\text{norm}}(P)$ of the entire generation distribution.
+where $P$ represents the original token-level probability distribution, and $\alpha H_{\text{norm}}(P)$ dynamically scales the adjustment based on model uncertainty.
 5 Experiments
@@
 EM F1 EM F1 EM F1 EM F1 EM F1 EM F1 EM F1 EM F1
-LLaMA2-7B Regular 44.74 54.71 55.28 68.01 54.21 59.24 39.90 52.61 32.93 45.46 38.85 50.59 36.03 36.62 43.13 52.46
+LLaMA2-7B Greedy 44.74 54.71 55.28 68.01 54.21 59.24 39.90 52.61 32.93 45.46 38.85 50.59 36.03 36.62 43.13 52.46
 CAD 44.13 54.49 55.26 68.04 54.14 59.21 38.35 51.12 31.74 43.70 38.14 48.58 36.11 36.69 42.55 51.69
 COIECD 42.03 51.48 57.04 70.06 57.03 63.14 40.93 54.78 34.40 48.48 38.79 51.69 34.98 35.63 43.60 53.61
-OURs 47.67 57.87 58.46 71.42 55.04 60.35 49.00 61.04 36.70 49.29 48.10 60.71 54.33 54.95 49.90 59.38
-LLaMA2-7B -Chat Regular 53.33 67.41 71.72 76.83 54.19 58.11 67.69 78.70 39.41 53.90 50.47 65.48 67.98 68.78 57.83 67.03
+OURs 47.35 57.43 58.12 70.97 54.35 59.70 48.02 60.02 36.51 49.06 47.74 60.23 53.12 53.63 49.32 58.72
+LLaMA2-7B -Chat Greedy 53.33 67.41 71.72 76.83 54.19 58.11 67.69 78.70 39.41 53.90 50.47 65.48 67.98 68.78 57.83 67.03
 CAD 52.86 67.16 71.70 76.83 54.16 58.11 65.89 77.91 38.46 53.26 48.89 65.00 68.04 68.85 57.14 66.73
 COIECD 53.14 67.03 72.26 77.33 55.04 58.98 68.32 79.56 40.00 54.55 52.39 66.84 69.48 70.13 58.66 67.77
-OURs 55.26 68.65 70.05 76.56 54.70 58.57 68.60 79.65 40.38 54.72 51.75 66.93 69.50 70.30 58.61 67.91
-LLaMA2-13B Regular 52.36 63.40 58.25 69.95 63.22 68.33 51.64 64.57 30.84 40.11 42.26 54.08 49.02 49.59 49.66 58.58
+OURs 55.31 68.61 69.21 75.95 54.25 58.13 68.49 79.76 40.53 54.81 51.69 66.92 69.50 70.30 58.43 67.78
+LLaMA2-13B Greedy 52.36 63.40 58.25 69.95 63.22 68.33 51.64 64.57 30.84 40.11 42.26 54.08 49.02 49.59 49.66 58.58
 CAD 51.53 63.11 58.25 69.92 63.13 68.32 49.94 63.44 29.94 39.04 41.40 52.27 49.07 49.63 49.04 57.96
 COIECD 50.21 60.96 59.19 71.49 65.97 71.00 52.73 65.93 35.66 50.58 42.35 54.37 48.35 48.84 50.64 60.45
-OURs 53.86 65.25 59.65 72.07 61.41 67.08 57.36 69.85 33.52 42.81 55.34 71.23 65.86 66.19 55.29 64.93
-LLaMA2-13B -Chat Regular 55.01 69.92 74.58 79.35 67.08 71.96 68.26 79.45 40.20 55.11 53.49 69.18 60.69 61.77 59.90 69.53
+OURs 53.13 64.21 59.65 72.07 61.41 67.08 56.54 68.56 33.26 42.20 55.34 71.23 65.86 66.19 55.03 64.51
+LLaMA2-13B -Chat Greedy 55.01 69.92 74.58 79.35 67.08 71.96 68.26 79.45 40.20 55.11 53.49 69.18 60.69 61.77 59.90 69.53
 CAD 54.44 69.66 74.58 79.37 67.01 71.95 66.77 78.70 39.44 54.44 52.89 68.63 60.83 61.92 59.42 69.24
 COIECD 56.15 70.43 73.87 78.96 67.28 71.93 68.49 80.39 40.75 56.16 53.69 69.81 62.47 63.21 60.39 70.13
-OURs 58.22 71.99 73.55 79.11 68.19 72.73 69.83 88.88 40.77 56.54 55.34 71.24 64.30 65.42 61.46 72.27
-Mistral-7B Regular 53.41 64.36 59.45 68.39 63.79 67.77 44.19 56.11 31.51 38.94 33.74 51.18 39.80 46.04 46.56 56.11
+OURs 57.76 71.69 73.04 78.77 68.19 72.73 69.66 80.76 40.78 56.24 55.36 71.31 64.03 65.20 61.26 70.96
+Mistral-7B Greedy 53.41 64.36 59.45 68.39 63.79 67.77 44.19 56.11 31.51 38.94 33.74 51.18 39.80 46.04 46.56 56.11
 CAD 41.57 56.01 57.88 67.48 63.64 68.65 34.08 47.55 25.78 35.37 23.18 41.46 26.96 35.97 39.01 50.36
 COIECD 46.43 58.32 44.30 51.77 54.82 59.17 50.50 60.98 40.05 52.78 42.12 56.58 59.53 61.89 48.25 57.36
-OURs 63.43 73.41 57.10 72.01 64.71 70.62 62.44 73.86 41.41 55.00 57.80 72.54 69.90 70.39 59.54 69.69
-Mistral-7B -Instruct Regular 58.70 72.18 69.64 75.61 44.42 49.63 67.28 79.37 39.79 54.72 52.29 66.93 66.90 67.83 57.00 66.61
+OURs 63.45 73.89 56.76 71.86 64.49 69.88 63.04 73.75 41.62 55.13 57.85 71.19 69.46 69.86 59.52 69.37
+Mistral-7B -Instruct Greedy 58.70 72.18 69.64 75.61 44.42 49.63 67.28 79.37 39.79 54.72 52.29 66.93 66.90 67.83 57.00 66.61
 CAD 49.30 64.81 70.23 75.95 45.42 50.96 59.97 72.92 34.97 51.90 42.63 58.49 52.04 53.75 50.65 61.25
 COIECD 59.74 72.59 64.92 72.15 37.09 42.66 68.45 81.03 40.84 55.96 53.54 68.72 72.81 73.81 56.77 66.70
-OURs 60.77 73.82 69.70 75.96 47.17 52.65 68.29 80.74 40.85 56.01 54.69 69.81 71.27 72.10 58.96 68.73
+OURs 60.55 73.49 69.70 75.96 47.17 52.65 68.30 80.62 40.85 56.23 54.78 69.72 71.38 72.12 58.96 68.68
 Table 1: Performance comparison of different decoding methods.
@@
 Bold indicates the best performance, and underlined indicates the second-best performance.
+Figure 5: Ablation Study for DAGCD. Left Part: Ablation 1 Detector Training Data Sizes, Center Part: Ablation 2 Top-Rank Constraint, Right Part: Ablation 3 Scaling Factor $\alpha$.
 5.1 Experimental Setup
-Models We conducted experiments on LLaMA and Mistral, including LLaMA2 (7B, 13B, 7B-Chat, 13B-chat) Touvron et al. (2023) and Mistral (7B, 7B-Instruct) Jiang et al. (2023).
-Datasets We conducted experiments on seven open-book question-answering (QA) datasets, representing a variety of QA tasks.
-These include multi-hop reasoning datasets (HotpotQA Yang et al.
-(2018)), long-form retrieval-based QA datasets (TriviaQA Joshi et al.
-(2017) and SearchQA Dunn et al.
-(2017)), single-paragraph extraction tasks (SQuAD Rajpurkar et al.
-(2016) and NewsQA Trischler et al.
-(2017)), and document-level QA datasets (NQ Kwiatkowski et al.
-(2019)).
-All datasets are formatted in the unified schema provided by the MrQA repository Fisch et al.
-(2019).
-Additionally, we used the artificially constructed NQ-swap dataset Longpre et al.
-(2021), designed to simulate conflicting or ambiguous scenarios by replacing entities.
-Detailed descriptions see Appendix F.3.
-Baselines We compare the proposed DAGCD method with three decoding strategies: Greedy Decoding, CAD Shi et al.
-(2024b), and COIECD Yuan et al.
-(2024).
-CAD and COIECD are specifically designed to mitigate context faithfulness hallucination.
-Implementation Details For DAGCD, the scaling factor $\alpha$ is set to 2 for pretrained models and 4 for instruction-tuned models to account for entropy variations.
+Datasets
+We evaluate context faithfulness on Open-Book Question-Answering (QA) datasets, where each question is paired with external context containing the correct answer.
+This setup ensures that only context-grounded answers are considered valid, allowing for the assessment of whether the model generates context-faithful hallucinations based on answer accuracy.
+Specifically, We conducted experiments on seven open-book QA datasets.
+For further details, refer to Appendix D.1.
+Implementation Details
+We used 100 samples constructed from a single dataset (HotpotQA) as the training set to train the Context Utilization Detector.
+The scaling factor $\alpha$ is set to 2 for pre-trained models and 4 for instruction-tuned models to account for entropy variations.
 The logistic regression classifier and utilization distribution are computed using the top-10 attention heads, with adjustments restricted to the top-10 ranked tokens ($U_{\text{top}}$).
-Further details on prompts, decoding strategies, and baseline configurations are provided in Appendix F.4 and F.5.
-Metrics Consistent with prior workJin et al. (2024a); Yuan et al. (2024); Wang et al. (2024), we use EM and F1 score metrics to evaluate the performance of the models on open-domain QA datasets.
+Further detailed settings see Appendix D.2.
+Metrics
+Consistent with prior workJin et al. (2024a); Yuan et al. (2024); Wang et al. (2024), we use EM and F1 score metrics to evaluate the performance of the models on open-book QA datasets.
+The LLMs used in our experiments and the baselines are detailed in Appendix D.3 and D.4.
 5.2 Model Performance Comparison
-Table 3 demonstrates the effectiveness of DAGCD across diverse QA datasets and model configurations.
-We summarize the findings based on datasets and model characteristics.
-Additionally, we evaluated DAGCD on summarization tasks, and the results demonstrate that DAGCD also achieves improvements (detailed results see Appendix F.6).
+Table 1 shows DAGCD’s effectiveness across diverse QA datasets and models. We also tested DAGCD on summarization tasks, yielding improvements. For details see Appendix D.5.
 5.2.1 Dataset-Level Observations
 DAGCD achieves consistent improvements across diverse QA tasks, including multi-hop reasoning, long-form retrieval, and document-level QA.
-HotpotQA, TriviaQA, SearchQA: DAGCD demonstrates strong performance in tasks requiring multi-paragraph reasoning and long-form retrieval.
-It achieves the highest gains on HotpotQA with a 10.02% EM and 9.05% F1 improvement on Mistral-7B.
-Additionally, DAGCD outperforms CAD and COIECD on TriviaQA and SearchQA, with significant improvements across different models.
-SQuAD, NewsQA, NQ: DAGCD demonstrates robust performance in single-paragraph and document-level tasks.
-On NQ, it achieves a 24.06% EM and 21.36% F1 increase on Mistral-7B compared to greedy decoding, while delivering consistent gains across SQuAD and NewsQA datasets.
-NQ-Swap: In adversarial scenarios simulated by NQ-Swap, DAGCD achieves significant improvements, including 30.10% EM and 24.35% F1 gains on Mistral-7B, underscoring its robustness in challenging conditions.
+HotpotQA, TriviaQA, SearchQA:
+DAGCD excels in multi-paragraph reasoning and long-form retrieval tasks.
+It achieves the highest gains on HotpotQA with a 18.80% EM and 14.81% F1 improvement on Mistral-7B.
+DAGCD also outperforms baselines on TriviaQA and SearchQA, showing significant improvements across models.
+SQuAD, NewsQA, NQ:
+DAGCD demonstrates robust performance in single-paragraph and document-level tasks.
+On NQ, it achieves a 71.46% EM and 39.10% F1 improvement on Mistral-7B over greedy decoding, while delivering consistent gains across SQuAD and NewsQA datasets.
+NQ-Swap:
+In adversarial scenarios simulated by NQ-Swap, DAGCD shows notable improvements, including 74.52% EM and 51.74% F1 gains on Mistral-7B, highlighting its robustness.
 5.2.2 Model-Level Observations
 DAGCD demonstrates broad applicability across different model families, sizes, and tuning variants.
-Model Families: DAGCD consistently improves performance across both LLaMA and Mistral families.
-On Mistral-7B, DAGCD achieves an average EM of 59.54 and an average F1 of 69.69, outperforming greedy decoding by 12.99% and 13.58%, respectively.
-LLaMA2 models also benefit significantly, on LLaMA2-7B, with average F1 improvements of 6.91% over greedy decoding.
-Model Sizes: DAGCD improves performance across models of varying model sizes, achieving an average EM and F1 increase of 6.77% and 6.91% on LLaMA2-7B, and 5.63% and 6.35% on LLaMA2-13B, respectively.
-Instruction-Tuned Models: DAGCD consistently achieves performance gains across instruction-tuned models, surpassing all baselines with the highest average EM and F1 scores.
-Figure 4: The impact of different $\alpha$ values on the performance of our method (fixed top-rank = 10).
-Figure 5: The impact of different Top-Rank constraint on the performance of DAGCD (fixed $\alpha=2$ and $4$).
+Model Families:
+DAGCD enhances performance across LLaMA and Mistral families. On Mistral-7B, DAGCD improves EM by 27.86% on Mistral-7B and 14.33% on LLaMA2-7B compared to greedy decoding.
+Model Sizes:
+DAGCD improves performance across models of varying model sizes, achieving an average EM and F1 increase of 14.33% and 11.93% on LLaMA2-7B, and 10.82% and 10.12% on LLaMA2-13B, respectively.
+Instruction-Tuned Models:
+Instruction-tuned models, after fine-tuning, show reduced uncertainty in generation probabilities, limiting our method’s improvement margin.
+However, DAGCD still surpasses all baselines with the highest performance.
 5.3 Ablation Study
-Impact of Different $\alpha$ We evaluated the impact of different values of $\alpha$ on model performance with the top-rank fixed at 10.
-The results, presented in Figure 4, indicate that $\alpha$ determines the adjustment intensity applied to the original generation distribution.
+We conducted three ablation experiments to evaluate the impact of variations in different modules on performance.
+Specifically, in each ablation experiment, the ablation module is adjusted while the other two modules are kept at their default settings.
+Ablation 1: Detector Training Data Sizes
+We tested the impact of detectors trained on different data sizes on actual inference performance.
+The results, shown in Figure 5 (left), demonstrate that our method consistently maintains strong performance across various training data sizes, achieving notable results even with just 100 training samples.
+Ablation 2: Top-Rank Constraint
+We evaluated various top-rank constraints on HotpotQA.
+Figure 5 (center), top-rank filtering reduces false positives, with F1 score initially improving as constraints loosen, then declining when overly relaxed.
+Ablation 3: Scaling Factor $\alpha$
+We evaluated the impact of different scaling factor $\alpha$ on model performance.
+The results, presented in Figure 5 (right), indicate that $\alpha$ determines the adjustment intensity applied to the original generation distribution.
 For pretrained models, optimal performance is achieved at $\alpha=2$, whereas for Chat models, the best performance is observed at $\alpha=4$.
-Impact of Different Top-Rank We evaluated F1 score variations under different top-rank filtering constraints on HotpotQA.
-As shown in Figure 5, top-rank filtering reduces false positives, with performance initially improving as constraints loosen, then declining when overly relaxed.
-Additional results and performance variations under different prompts in Appendix F.7 and F.8.
+Additional results and performance variations under different prompt templates see Appendix E.
 6 Discussion and Analysis
@@
 (2024b) and COIECD Yuan et al.
-(2024), improve contextual alignment but often rely on multi-step processes, resulting in significant computational overhead.
-In contrast, our dynamic decoding approach integrates context adjustments directly into the generation process, providing both efficiency and real-time optimization.
-Lightweight Context Utilization Detector Using a logistic regression-based Context Utilization Detector, our method enables real-time adjustments with minimal computational overhead.
-This detector is efficient compared to resource-intensive alternatives, such as integrated gradients or exhaustive attention head manipulation.
-Single-Pass Decoding with Real-Time Faithfulness Optimization By integrating the Context Utilization Detector directly into the decoding process, our method eliminates redundant steps such as output comparisons or external consistency checks.
+(2024), improve contextual alignment but rely on multi-step processes, causing significant computational overhead.
+In contrast, DAGCD incorporates context adjustments during generation, ensuring both efficiency and real-time optimization.
+Lightweight Context Utilization Detector Using a logistic regression-based Context Utilization Detector, our method enables real-time adjustments with minimal computational cost.
+This detector is more efficient than resource-heavy methods like integrated gradients or attention head manipulation.
+Single-Pass Decoding with Real-Time Faithfulness Optimization By integrating the Context Utilization Detector directly into the decoding process, our method removes redundant steps like output comparisons or external consistency checks.
 During generation, attention-based context utilization signals are leveraged in real time to proactively enhance faithfulness.
-This single-step strategy ensures that the output remains aligned with the input context without requiring additional post-processing, while maintaining the theoretical time complexity of regular decoding.
+This single-step strategy ensures that the output aligns with the input context without extra post-processing, while maintaining the theoretical time complexity of greedy decoding.
 6.2 Interpretability Through Attention: Insights into Context Utilization
-By systematically analyzing attention mechanisms, our approach uncovers how retrieved context influences the generation process and provides interpretable insights into the model’s behavior.
-Feature-Based Attention Analysis: Using naturally occurring cases (e.g., failure in closed-book settings but success in open-book settings), we systematically isolate attention patterns that are indicative of context utilization.
-A logistic regression model trained on these patterns identifies the most relevant attention heads with high accuracy, providing a quantitative measure of each head’s contribution to context alignment.
-Transparent Decision-Making: The feature coefficients of the logistic regression model directly map to the importance of specific attention heads.
-This transparency allows for intuitive interpretation, making it clear which heads are most responsible for leveraging context tokens during generation.
+By systematically analyzing attention mechanisms, our approach uncovers how retrieved context influences the generation process and provides interpretable insights into the behavior of the model.
+Feature-Based Attention Analysis Using natural cases, such as failure in closed-book settings but success in open-book settings, we isolate attention patterns that are indicative of context utilization.
+A logistic regression classifier trained on these patterns identifies the relevant attention heads with high accuracy, quantifying their contributions.
+Transparent Decision-Making The feature coefficients of the logistic regression model directly map to the importance of specific attention heads.
+This transparency allows for intuitive interpretation, clarifying which heads are most responsible for leveraging context tokens during generation.
 7 Related Work
@@
 Lei et al.
-(2023) proposed a post-generation editing strategy that uses natural language inference to classify and revise hallucinated segments in generated content.
+(2023) proposed a post-generation editing strategy using natural language inference to classify and revise hallucinated segments.
 Choi et al.
 Choi et al.
-(2023) introduced Knowledge-Constrained Decoding, which detects hallucinations at the token level during generation and reweights token distributions to guide output.
-Similarly, Chuang et al.
+(2023) introduced Knowledge-Constrained Decoding, detecting hallucinations during generation and reweighting token distributions to guide output.
 Chuang et al.
-(2024a) proposed Lookback Lens Guided Decoding, which selects the most faithful output among multiple candidates to improve consistency.
+Chuang et al.
+(2024a) proposed Lookback Lens Guided Decoding, selecting the most faithful output among candidates to improve consistency.
 For mitigation, CAD Shi et al.
@@
 While effective, most existing methods face limitations such as high computational overhead and reliance on multiple decoding passes.
-In this work, we propose a real-time solution that integrates context utilization signals directly into the decoding process, achieving efficient and faithful generation without additional processing steps.
-7.2 Attention Mechanisms and Interpretability
+In this work, we propose a real-time solution that integrates context utilization signals directly into the decoding process, achieving efficient and faithful generation, without the need for additional processing steps.
+7.2 Attention and Interpretability
 The attention mechanism provides valuable insights into how models prioritize different parts of an input sequence Clark et al.
-(2019a); Geva et al.
+(2019); Geva et al.
 (2023) and has become central to understanding Transformer-based LLMs Vashishth et al.
@@
 In LLMs, attention heads often perform distinct roles, such as capturing syntactic dependencies or aligning semantic relationships Olsson et al.
-(2022); Zheng et al.
+(2022b); Zheng et al.
 (2024); Jin et al.
@@
 (2024a) further investigate the contributions of attention scores and neurons to model outputs.
-Unlike prior work that focuses on individual heads or neurons, our study examines the collaborative patterns among multiple attention heads.
-By analyzing how attention mechanisms collectively utilize contextual tokens, we provide a holistic view of their role in aligning outputs with user-provided context, offering new insights into Transformer-based model behavior.
+Unlike previous work, which focused on individual heads, our study examines the collaborative patterns of multiple attention heads.
+By analyzing how attention mechanisms collectively utilize contextual tokens, we provide a holistic view of their role in aligning outputs with user-provided context.
 8 Conclusion
-In this paper, we addresses context faithfulness hallucinations in LLMs by proposing Dynamic Attention-Guided Context Decoding (DAGCD), a lightweight framework that integrates attention distributions and entropy-based uncertainty signals to amplify contextually relevant tokens during generation.
-Our analysis revealed a strong correlation between high uncertainty and hallucinations, and probing experiments validated that attention mechanisms encode signals indicative of contextual utilization.
-Experiments across multiple QA datasets demonstrated that DAGCD achieves consistent improvements in context faithfulness, robustness, and scalability, providing an effective solution for context-sensitive generation tasks.
+In this paper, we mitigate context faithfulness hallucinations in LLMs by proposing Dynamic Attention-Guided Context Decoding, a lightweight framework that integrates attention distributions and entropy-based uncertainty signals to amplify contextually relevant tokens during generation.
+Our analysis revealed a strong correlation between high uncertainty and hallucinations, and probing experiments validated that attention mechanisms encode signals indicative of contextual utilization, and further demonstrated that this signal is a fundamental mechanism in Transformer-based LLMs.
+Experiments across multiple open-book QA datasets demonstrated that DAGCD achieves consistent improvements in context faithfulness, robustness, and scalability, providing an effective solution for context-sensitive generation tasks.
 Limitations
@@
 Extending the method to diverse domains and tasks could further validate its robustness and scalability.
-Appendix A Prediction Uncertainty Analysis Details
-A.1 Dataset and Sampling
-We selected 6,000 samples from the MrQA training set Fisch et al.
-(2019), distributed evenly across six sub-datasets.
-Each sub-dataset contained 500 correctly answered samples and 500 incorrectly answered samples, evaluated using the Exact Match (EM) metric.
-A.2 Prompt and Decoding Settings
-The Prompt 1 template Yuan et al.
-(2024), as shown in Appendix G, was used for response generation.
-LLaMA2-7B Touvron et al.
-(2023) served as the backbone model, with greedy decoding ensuring deterministic outputs.
+Appendix A Details of Experiment "2.1 Uncertainty Leads to Unfaithful Answers"
+A.1 Dataset Used for Analysis
+We randomly sampled 6,000 instances from the MrQA training set Fisch et al.
+(2019), which consolidates six open-domain question answering datasets.
+Specifically, we selected 500 correctly answered and 500 incorrectly answered instances from each sub-dataset.
+An answer was considered correct if it achieved an Exact Match (EM) with the reference answer.
+A.2 Model Used for Analysis, Prompts, and Decoding Method
+We conduct our analysis using LLaMA2-7B Touvron et al. (2023) as the target model. The answers are generated using the following prompt:
+"Given the following information: {context} Answer the following question based on the given information with one or a few words: {question} Answer:"
+To ensure deterministic outputs, we employ greedy decoding.
 A.3 Computation Process
-For each sample, we calculated two metrics to quantify uncertainty.
-First, the normalized entropy, defined in Equation 1, measures the dispersion of probabilities across the vocabulary, providing an overall view of the model’s uncertainty.
-Second, the top-1 token probability represents the likelihood of the most probable token, offering a complementary perspective on the model’s prediction confidence.
-These metrics focus on the initial generated token to analyze how uncertainty affects response faithfulness.
-A.4 Dataset and Sampling
-We selected 6,000 samples from the training set Fisch et al.
-(2019) for this analysis (The dataset consists of both the question and the context containing gold answer.).
-These samples were evenly distributed across six sub-datasets, with 1,000 samples drawn from each.
-For each sub-dataset, consisting of 500 correctly answered samples and 500 incorrectly answered samples (given context), the correctness of the responses is evaluated using the Exact Match (EM) metric.
-A.5 Uncertainty Metrics Calculation Settings
-We employed the Prompt 1 template Yuan et al.
-(2024) (as shown in Appendix G) for generation with LLaMA2-7B Touvron et al.
-(2023), using greedy decoding to ensure deterministic predictions.
-For each sample, we computed the normalized entropy of the distribution when the model generates the first answer token, while simultaneously recording the probability of the top-1 token in the distribution.
-Appendix B Attention Signal Analysis Details
-B.1 Dataset Construction
-We constructed the dataset by selecting samples from the MrQA training set dataset Fisch et al.
-(2019), focusing on cases where the model’s output changed from incorrect to correct after context concatenation.
-A total of 2,000 positive samples (utilized tokens) and 2,000 negative samples (non-utilized tokens) were selected.
-And each sample consists of a label (utilized or non-utilized token) and the corresponding attention ratio feature vector $\mathbf{v}_{j}$ (under Prompt 1).
-The dataset was evenly split into training (50%) and testing (50%) sets.
-B.2 Training and Evaluation
+For each sample, we calculated two metrics to quantify uncertainty and confidence.
+First, the Normalized Entropy (NE) measures the dispersion of probabilities across the vocabulary, providing an overall view of the model’s uncertainty.
+And NE is defined as:
+$H_{\text{norm}}(P)=-\frac{\sum_{i=1}^{N}P_{i}\log P_{i}}{\log N},$ (7)
+where $P$ represents the token-level probability distribution, and $N$ denotes the vocabulary size.
+Second, the Maximum Softmax Probability (MSP) represents the likelihood of the most probable token, offering a complementary perspective on the model’s prediction confidence.
+These metrics focus on the initial generated token distribution to analyze how uncertainty affects response faithfulness.
+A.4 Spearman Correlation Analysis
+To further investigate the relationship between uncertainty and model errors (i.e., the inability to faithfully respond to the input context), we conducted a Spearman correlation analysis.
+Specifically, we used the normalized entropy of the token level probability distribution to measure the model’s token-level uncertainty, and evaluated answer accuracy using the F1 score.
+We then analyzed the correlation between answer accuracy and uncertainty during answer generation, with results presented in Table 2.
+Our findings reveal a significant negative correlation, which becomes more pronounced after concatenating the context—i.e., higher uncertainty corresponds to lower answer accuracy (in open-book QA tasks, lower accuracy indicates that the model’s response deviates from the provided context).
+This analysis further suggests that token-level uncertainty is strongly associated with unfaithful answers.
+Model w/o context w/ context
+LLaMA2-7B -0.43 -0.53
+LLaMA2-7B-Chat -0.27 -0.33
+LLaMA2-13B -0.30 -0.51
+LLaMA2-13B-Chat -0.26 -0.32
+Mistral-7B -0.23 -0.22
+Mistral-7B-Instruct -0.30 -0.33
+Table 2: Spearman correlation analysis.
+We examine the relationship between F1 scores (pred-ans vs.
+gold-ans) and norm-entropy (generation distribution) under both w/o and w/ context settings ($p\ll 0.05$ for all models).
+A.5 Probability Gap Between The Golden Answer Token and The Ranked Top-1 Token
+As shown in Figure 6, we analyze the wrong answer samples from A.1 by calculating the average probability gap between the gold answer token and the highest-probability token.
+Figure 6: Probability Gap. For incorrect responses, the probability gap between the gold answer token and the highest-probability token.
+Appendix B Details of "3 Context Utilization Signal in Attention"
+B.1 LR Classifier Training and Evaluation
 We trained a Logistic Regression (LR) classifier using 5-fold cross-validation.
@@
   • Mistral: 7B, 7B-Instruct
-Appendix C Details of LR Classifier Generalization and Feature Analysis
-C.1 Cross-Prompt Test Results
+B.2 Cross Prompt Templates Testing
 To evaluate the robustness of the classifier under different prompts, we reconstructed the attention ratio feature vectors using Prompts 2, 3, and 4.
@@
 The classifier, trained using Prompt 1, was then tested on these alternative prompts.
+(Templates shown in Figure 15)
 The results, shown in Table 3, demonstrate that the classifier maintains an ACC exceeding 97% and AUC above 0.99 across all prompts.
@@
 Figure 9: Heatmap of feature coefficients for LR. The heatmap of feature coefficients for LR classifiers trained using attention ratios from different LLMs as features.
-C.2 Analysis of the Importance of Different Features
+Figure 10: Performance distribution of LR with single features.
+An LR classifier is trained using the attention ratio from a single head as features.
+The figure shows the ACC distribution of LR classifiers trained with attention ratios from different heads on the test set.
+Figure 11: Performance of the LR classifier trained with Bottom-$K$ features.
+Based on the ACC of LR classifiers trained using the attention ratio from a single head as features, the Bottom-$K$ heads with the lowest ACC are selected.
+The LR classifier is then retrained using the attention ratios of these Bottom-$K$ heads as features, and its performance on the test set is presented.
+Appendix C Analyzing Context Utilization Signal Contribution Across Heads
+To further explore the role of attention heads in encoding contextual utilization signals, we conducted a detailed analysis to examine the strength and distribution of signals across heads.
+This section presents insights into the concentration of strong signals in certain heads, the independent utility of individual heads, and the complementarity of weaker heads.
+C.1 Analysis of the Importance of Different Features
 In analyzing the LR classifier trained with attention values from all attention heads as features, we found that most feature coefficients had small absolute values.
@@
 Therefore, all subsequent inference employs LR classifiers trained with the Top-10 features.
-Figure 9: Performance distribution of LR with single features.
-An LR classifier is trained using the attention ratio from a single head as features.
-The figure shows the ACC distribution of LR classifiers trained with attention ratios from different heads on the test set.
-Figure 10: Performance of the LR classifier trained with Bottom-$K$ features.
-Based on the ACC of LR classifiers trained using the attention ratio from a single head as features, the Bottom-$K$ heads with the lowest ACC are selected.
-The LR classifier is then retrained using the attention ratios of these Bottom-$K$ heads as features, and its performance on the test set is presented.
-Appendix D Analyzing Context Utilization Signal Contribution Across Heads
-To further explore the role of attention heads in encoding contextual utilization signals, we conducted a detailed analysis to examine the strength and distribution of signals across heads.
-This section presents insights into the concentration of strong signals in certain heads, the independent utility of individual heads, and the complementarity of weaker heads.
-D.1 Signal Strength and Concentration in Heads
+C.2 Signal Strength and Concentration in Heads
 To identify influential heads, we visualized the coefficients of the trained Logistic Regression (LR) classifier, which were derived from the attention ratio features of all heads.
@@
 This indicates the presence of more high-performing heads in the Mistral model, which may explain why our method achieves greater improvements on Mistral compared to other models.
-D.2 Complementary Contributions of Weaker Heads
+C.3 Complementary Contributions of Weaker Heads
 Although most attention heads have limited standalone utility, we observe that combining weaker heads into subsets significantly improves classification performance.
@@
 This finding highlights the complementarity of weaker heads, as their aggregated signals collectively achieve robust token classification.
-D.3 Data-Dependent and Generalizable Heads
-To investigate whether the influential heads identified in one dataset maintain their significance across other datasets, we conducted cross-dataset experiments.
-For datasets $A$, $B$, and $C$, we analyzed the top 10 heads ranked by LR coefficients and observed significant differences in head rankings across datasets.
-However, when considering the intersection of the top 100 heads across all three datasets, we identified a subset of heads that consistently exhibited high importance.
-Using this subset of generalizable heads, we retrained the classifier on dataset $A$ and tested it on datasets $B$ and $C$.
-The model achieved high performance across all datasets, suggesting the presence of data-independent heads that encode universal contextual utilization signals.
-Conversely, dataset-specific top heads exhibited superior performance on their respective datasets, indicating a degree of data dependency in some heads’ contributions.
 Summary of Findings
@@
     Complementarity: Weaker heads collectively provide complementary signals, enabling robust classification when aggregated.
-  3. 3.
-    Data Dependency and Generalization: Some heads exhibit data-specific importance, while others demonstrate generalizability across datasets, reflecting a balance between dataset-specific and universal contributions.
 These findings highlight the nuanced roles of attention heads in contextual token utilization and provide a foundation for further exploration of their properties and applications.
-Appendix E Models and Datasets
-E.1 Model Details
-The LLMs used in this work, along with its HuggingFace ID, is as follows:
-  • LLaMA2-7B: meta-llama/Llama-2-7b-hf
-  • LLaMA2-7B-Chat: meta-llama/Llama-2-7b-chat-hf
-  • LLaMA2-13B: meta-llama/Llama-2-13b-hf
-  • LLaMA2-13B-Chat: meta-llama/Llama-2-13b-chat-hf
-  • Mistral-7B: mistralai/Mistral-7B-v0.1
-  • Mistral-7B-Instruct: mistralai/Mistral-7B-Instruct-v0.1
-Appendix F Experimental Details
+Appendix D Experimental Details
+D.1 Dataset Details
+We conducted experiments on seven open-book question-answering (QA) datasets, representing a variety of QA tasks.
+These include multi-hop reasoning datasets (HotpotQA Yang et al.
+(2018)), long-form retrieval-based QA datasets (TriviaQA Joshi et al.
+(2017) and SearchQA Dunn et al.
+(2017)), single-paragraph extraction tasks (SQuAD Rajpurkar et al.
+(2016) and NewsQA Trischler et al.
+(2017)), and document-level QA datasets (NQ Kwiatkowski et al.
+(2019)).
+All datasets are formatted in the unified schema provided by the MrQA repository Fisch et al.
+(2019).
+Additionally, we used the artificially constructed NQ-swap dataset Longpre et al.
+(2021), designed to simulate conflicting or ambiguous scenarios by replacing entities.
+D.1.1 Dataset Categories and Statistics
 The datasets used in this study include seven open-book question-answering (QA) datasets, grouped into three categories based on their QA task characteristics: multi-hop reasoning, long-form retrieval-based QA, and single-paragraph extraction tasks.
@@
 Detailed descriptions and dataset statistics are provided below.
-F.1 Dataset Categories and Statistics
 Multi-Hop Reasoning (HotpotQA).
@@
 These datasets test the model’s ability to focus on relevant content in lengthy contexts and generate precise answers.
-Single-Paragraph Extraction (SQuAD).
+Single-Paragraph Extraction (SQuAD, NewsQA).
 SQuAD Rajpurkar et al.
@@
 It evaluates the model’s ability to resolve conflicts and faithfully utilize context.
-F.2 Dataset Sources and Formats
+D.1.2 Dataset Sources and Formats
 All datasets are standardized in the unified schema provided by the MrQA repository (Huggingface ID: mrqa-workshop/mrqa), except for NQ-Swap, which is sourced from a separate repository (Huggingface ID: pminervini/NQ-Swap).
@@
 All datasets have been preprocessed to ensure compatibility with our experimental framework.
-F.3 Dataset Statistics
+D.1.3 Dataset Statistics
 Table 4 presents the size of the datasets used in this study to evaluate model performance.
@@
 Table 4: Dataset statistics. A summary of the dataset sizes used for evaluation across different datasets.
-F.4 Baseline Method Configurations
+D.2 Implementation Details
+[…]
+D.3 Model Details
+The LLMs used in this work, along with its HuggingFace ID, is as follows:
+  • LLaMA2-7B: meta-llama/Llama-2-7b-hf
+  • LLaMA2-7B-Chat: meta-llama/Llama-2-7b-chat-hf
+  • LLaMA2-13B: meta-llama/Llama-2-13b-hf
+  • LLaMA2-13B-Chat: meta-llama/Llama-2-13b-chat-hf
+  • Mistral-7B: mistralai/Mistral-7B-v0.1
+  • Mistral-7B-Instruct: mistralai/Mistral-7B-Instruct-v0.1
+D.4 Baseline Configurations
+We compare the proposed method DAGCD with three decoding strategies: Greedy Decoding, CAD Shi et al.
+(2024b), and COIECD Yuan et al.
+(2024).
+CAD and COIECD are specifically designed to mitigate context faithfulness hallucination.
 We implemented the baseline methods with their recommended hyperparameter settings for fair comparisons:
@@
   • COIECD Yuan et al. (2024): The entropy regularization parameter $\lambda$ was set to 0.25, and the contrastive adjustment factor $\alpha$ was set to 1.
-F.5 Implementation Details
-All experiments utilized a unified prompt template (Prompt 1) to ensure consistency across methods.
-The prompt format is detailed in Appendix G.
-For decoding, greedy decoding was employed to produce deterministic outputs and facilitate direct comparisons across methods.
-All models run on NVIDIA A100 GPUs.
 Model Decoding ROUGE-L factKB BERT-P BERT-R BERT-F1
-LLaMA2-7B Regular 0.2081 0.9932 0.9000 0.7997 0.8465
+LLaMA2-7B Greedy 0.2081 0.9932 0.9000 0.7997 0.8465
 CAD 0.2361 0.9786 0.9054 0.8016 0.8514
@@
 OURs 0.2134 0.9856 0.9210 0.8026 0.8576
-LLaMA2-7B-Chat Regular 0.2368 0.9846 0.9056 0.8035 0.8515
+LLaMA2-7B-Chat Greedy 0.2368 0.9846 0.9056 0.8035 0.8515
 CAD 0.2082 0.9417 0.9001 0.7997 0.8466
@@
 Bold denotes the best performance, while underlined indicates the second-best performance.
-F.6 Results on Summarization Tasks
+D.5 Results on Summarization Tasks
 To validate the performance of our approach on long-form answer generation tasks, we conducted experimental evaluations on the CNN_DM See et al.
@@
 The experimental results, as shown in Table 5, demonstrate that our method achieves significant improvements on both the pretrained and chat versions of LLaMA2.
-F.7 Additional Results of The Impact of Different $\alpha$
+Appendix E Detailed Results of "5.3 Ablation Study"
+E.1 Additional Results of "Ablation 1: Detector Training Data Sizes"
+Table 12 shows the performance variations of all LLMs used in this study when trained with detectors on different amounts of data.
+As observed, our method maintains consistent performance across various data sizes for all LLMs.
+Figure 12: Detector Training Data Size Validation. The variation in inference performance across different models when using detectors trained on varying amounts of data.
+E.2 Additional Results of "Ablation 3: Scaling Factor $\alpha$"
 We additionally evaluated the performance variations of our method on Mistral-7B and Mistral-7B-Instruct under different scaling factors $\alpha$.
@@
 This indicates that different models may require different optimal scaling factors for the best performance.
-F.8 Impact of Different Prompts
+Figure 13: The performance on HotpotQA for DAGCD under different scaling factors.
+E.3 Impact of Different Prompts
 To assess robustness to prompt variations, we tested multiple prompts from prior studies Zhou et al.
@@
 The results show that DAGCD consistently outperforms baselines across all tested prompts, demonstrating its adaptability to diverse input formats and reliability across QA tasks.
-Figure 11: The performance on HotpotQA for DAGCD under different scaling factors.
 Figure 14: Performance variations across different prompt templates.
 The figure shows F1 score variations on the HotpotQA and NewsQA datasets for Greedy Decoding and OURs (DAGCD) under different prompt templates.
-Appendix G Prompt Templates
+Appendix F Prompt Templates
 Figure 15 shows the prompt templates used in this paper.
 Figure 15: Prompt templates used in this paper.
-Appendix H Symbol Explanation
-The symbols and corresponding meanings defined in this paper, as shown in Table 7.
-Symbol Meaning
-$H_{\text{norm}}(P)$ Normalized entropy of distribution $P$
-$v_{j}$ Attention ratio feature vector for token $j$
-$v_{j}^{(K)}$ Top-K attention ratio feature vector for token $j$
-$a_{l,h}^{j}$ The raw attention of the $h$-th head in the $l$-th layer for token $j$
-$r_{l,h}^{j}$ The attention ratio of the $h$-th head in the $l$-th layer for token $j$
-$s_{j}$ The utilization score for token $j$
-$c_{k}$ The coefficient of the $k$-th feature in the LR classifier
-$w_{k}$ The normalized coefficient of the $k$-th feature in the LR classifier
-$U$ The utilization distribution
-$u_{i}$ The normalized utilization score for token $j$
-$U_{top}$ The subset of $U$ for tokens ranked in the top-$R$ of the generation distribution
-Table 7: Annotation table.